### v2

# Rethinking Multi-Modal Learning from Gradient Uncertainty

Abstract
Multi-Modal Learning (MML) integrates information from diverse modalities to improve predictive accuracy. While existing optimization strategies have made significant strides by mitigating gradient direction conflicts, we revisit MML from a gradient-based perspective to explore further improvements. Empirically, we observe an interesting phenomenon: performance fluctuations can persist in both conflict and non-conflict settings. Based on this, we argue that: beyond gradient direction, the intrinsic reliability of gradients acts as a decisive factor in optimization, necessitating the explicit modeling of gradient uncertainty. Guided by this insight, we propose Bayesian-Oriented Gradient Calibration for MML (BOGC-MML). Our approach explicitly models gradients as probability distributions to capture uncertainty, interpreting their precision as evidence within the framework of subjective logic and evidence theory. By subsequently aggregating these signals using a reduced Dempster’s combination rule, BOGC-MML adaptively weights gradients based on their reliability to generate a calibrated update. Extensive experiments demonstrate the effectiveness and advantages of the proposed method.
1 Introduction
Humans perceive the world through multiple sensory modalities, e.g., vision, hearing, and touch. This biological mechanism inspires the development of Multi-Modal Learning (MML), which aims to learn a model by integrating information from different modalities to achieve great performance (Huang et al., 2021; Wang et al., 2023b; Yuan et al., 2024; Song et al., 2025; Ma et al., 2025). Many existing approaches construct multiple unimodal losses together with a multi-modal fusion loss, and then linearly combine them for backpropagation (Yang et al., 2025; Huang et al., 2022) (Figure 2). These frameworks have been successfully applied across various fields, e.g., healthcare (Muhammad et al., 2021; Shi et al., 2025), autonomous systems (Agbley et al., 2021; Feng et al., 2020), human-computer interaction (D’mello and Kory, 2015; Li et al., 2020), and sentiment analysis (Poria et al., 2017, 2015; Zhang et al., 2024a).
Figure 1: Illustration of different gradient aggregation.
From an optimization perspective of MML, each update is jointly driven by two gradient sources: (i) unimodal gradients derived from modality-specific objectives; and (ii) multi-modal gradients induced by fusion, alignment, or aggregation objectives (Figure 2b). To enable efficient optimization, many approaches focus on the gradient directions in MML (Yu et al., 2020; Huang et al., 2022; Chen et al., 2018). Prior works show that gradients from different objectives can be in conflict (Peng et al., 2022; Zhang et al., 2024b), and has proposed a series of methods to mitigate such conflicts (Wei et al., 2024; Wei and Hu, 2024; Wang et al., 2023c; Yue et al., 2025). However, through empirical analysis, we observe an interesting phenomenon (Figure 3): in both conflict and non-conflict settings, while mitigating gradient conflicts generally boosts performance, the model can still underperform in certain cases.
Figure 2: Illustration of the MML training framework. (a) existing loss-level fusion paradigm, where unimodal losses and a multi-modal fusion loss are linearly combined. (b) gradient aggregation with fixed weights of existing methods vs. our uncertainty-driven fusion that adaptively reweights gradients based on their per-dimension reliability. (c) the overall MML pipeline.
Based on the above phenomenon, we argue that gradient directions are crucial for performance, yet the reliability of gradients themselves also plays an important role in MML. We propose to analyze this from the perspective of gradient uncertainty. A gradient is a high-dimensional vector that evolves during training, with each dimension being the partial derivative w.r.t. a model parameter (Goodfellow et al., 2016). Unlike prior methods, we treat gradients as random variables and characterize them with probability distributions. Gradient uncertainty can have an important influence in gradient aggregation. Take the two-dimensional update as an example (Figure 1), the unimodal gradient (yellow) has lower uncertainty along the $x$-axis but higher uncertainty along the $y$-axis, suggesting that it provides a more certain signal for updating toward the negative $x$ direction while being more uncertain for updating toward the positive $y$ direction. Similarly, the multi-modal gradient (blue) has lower uncertainty along the $y$-axis but higher uncertainty along the $x$-axis, suggesting that it is more certain for updating toward the negative $y$ direction while being more uncertain for updating toward the positive $x$ direction. In this case, both naive additive fusion and fusion after resolving conflicts can yield an update that deviates from the more certain components of the two gradients. This indicate that the more certain gradient components may be disturbed by the more uncertain ones, leading to suboptimal or even harmful update directions and thus limiting model performance. It also explains why performance degradation still occur under a non-conflict regime. Moreover, our experiments in Section 5 further show that explicitly accounting for gradient uncertainty during optimization can yield additional performance gains, highlighting that gradient reliability also plays a critical role in MML optimization. Thus, both gradient direction and gradient reliability are crucial for MML performance. Motivated by this, we aim to incorporate gradient uncertainty into the MML optimization to potentially mitigate the performance degradation.
To this end, we propose a Bayesian-Oriented Gradient Calibration method for MML (BOGC-MML), which explicitly models the uncertainty of each gradient to achieve more effective optimization. Instead of treating the gradients as deterministic values, we model them as random variables represented by probability distributions. Specifically, inspired by Bayesian inference (Box and Tiao, 2011; Dempster, 1968), we first model the gradients as probability distributions (Figure 2). Next, drawing from the “belief-uncertainty” in subjective logic (Jsang, 2018) and evidence theory (Bao et al., 2021; Sensoy et al., 2018), we propose an efficient mapping function (Theorem 4.1) that transforms the precision of the gradient distributions into a set of evidence. Furthermore, based on the relationship between evidence, belief, and uncertainty in subjective logic, we derive the following during the gradient aggregation process: (i) belief values, representing the confidence level of the gradients at each dimension; (ii) overall uncertainty of the gradient, reflecting the reliability of this gradient as an update signal during optimization. Finally, we use a reduced Dempster’s combination rule (Han et al., 2022) to aggregate gradients, yielding a calibrated update direction that accounts for both belief and uncertainty. It adaptively weights gradients to amplify low-uncertainty dimensions and down-weight high-uncertainty ones, thus achieve more effective optimization. Extensive experiments demonstrate its effectiveness.
The main contributions can be summarized as follows: (i) We revisit MML from an optimization perspective and argue that the reliability of gradients is also a critical factor for MML through empirical analysis that model can still underperform even mitigating gradient conflicts. (ii) We propose BOGC-MML, a novel Bayesian-oriented gradient calibration method for MML. It explicitly models the uncertainty of gradients and incorporates the dimensional differences between gradients during optimization, making the model update towards the optimal direction. (iii) We present an effective method for estimating evidence from probability distributions with theoretical guarantees. It maps the precision of a probability distribution to its corresponding evidence, enabling more effective gradient aggregation. (iv) Extensive experiments across various benchmarks demonstrate the effectiveness and validity of our proposed method, showcasing its advantages in practice.
(a) Conflict Scenario.
(b) Non-conflict Scenario.
Figure 3: Results of the discovery experiment.
2 Related Work
Multi-modal learning (MML) aims to integrate complementary signals from modalities such as images, text, and audio to improve performance (Baltrušaitis et al., 2018; Ngiam et al., 2011). A line of work tackles fusion from an optimization viewpoint by aggregating modality gradients and adaptively adjusting each modality’s contribution during training (Peng et al., 2022; Tsai et al., 2019; Wang et al., 2024, 2020; Peng et al., 2022; Fan et al., 2023; Wei and Hu, 2024). For example, (Wei and Hu, 2024) resolves conflicts between unimodal and multi-modal objectives via Pareto-based gradient integration; (Yue et al., 2025) uses an evidential dissonance measure to assess view quality and performs dissonance-aware belief integration with an evidential gradient penalty for robust multi-view fusion; (Jiang et al., 2025) modifies gradients via a flat gradient modification strategy (with SAM-style smoothing) to enhance cross-modal interaction and alleviate modality imbalance. However, these methods often rely on fixed coefficients and implicitly treat gradients as equally reliable, overlooking gradient uncertainty and dimension-wise sensitivity. We address this by explicitly modeling each modality gradient as a distribution and steering fusion toward more reliable update directions. More discussions are provided in Appendix C.
3 Problem Settings and Analyses
In this section, we first formalize important concepts and theories in this work. Subsequently, we investigate the impact of gradient reliability on model performance through empirical analysis, which also motivates our proposed method.
3.1 Problem Settings
In this subsection, we introduce the problem settings and related theories of this work. For more details and illustration of relevant concepts and theories, e.g., subjective logic and evidence theory, dempster-shafer evidence theory, etc., are provided in Appendix C.1.
Multi-Modal Learning In MML, the goal is to make accurate predictions by effectively integrating information from multiple modalities. To achieve this, it is common to jointly optimize via multi-modal fusion loss and unimodal losses (Wang et al., 2020) (Figure 2a). The former is computed based on the combined representation of all modalities and directly guides the model’s predictions; the latter supervises the learning of each modality representation, encouraging the model to extract important features from each modality.
Given a dataset $D^{i}=\{(x_{t}^{i},y_{t})\}_{t=1}^{n}$ where $x_{t}^{i}$ represents the input of the $i$-th modality for the $t$-th sample, and $y_{t}$ is the corresponding label. The overall loss function is defined as:
$\mathcal{L}_{\rm total}=\mathcal{L}_{\rm fusion}+\varphi{\textstyle\sum_{i=1}^{M}}\mathcal{L}_{\rm uni}^{i},$ (1)
where $\mathcal{L}_{\rm fusion}$ is the multi-modal fusion loss, $\mathcal{L}_{\rm uni}^{i}$ is the unimodal loss for the $i$-th modality, $\varphi$ is a regularization parameter, and $M$ is the total number of modalities.
3.2 Empirical Evidence and Motivation
Existing MML methods typically sum unimodal gradients and multi-modal gradients or mix them with fixed ratios during optimization (Yu et al., 2020; Wang et al., 2020; Huang et al., 2022; Wei and Hu, 2024; Wei et al., 2024). However, each update of unimodal encoders is jointly driven by two gradient sources: (i) unimodal gradients from unimodal losses, and (ii) multi-modal gradients from fusion. Both sources usually fluctuate across mini-batches and training stages, and a fixed fusion rule may not guarantee consistently beneficial updates, affecting performance. We further conduct an experiment to investigate this phenomenon.
Specifically, we adopt the Uniform baseline as the reference without any additional gradient aggregation strategy, and use G-Blending (Wang et al., 2020) as a representative method that fuses unimodal gradients with multi-modal fusion gradients during updates. We partition training scenarios into non-conflict and conflict cases and evaluate the final performance separately for each case. We further repeat G-Blending with the same configuration and report Worst Acc on CREMA-D (Cao et al., 2014). As shown in Figure 3(b), under the non-conflict setting, G-Blending slightly outperforms Uniform in most repeats (green bars above the yellow bar), but still underperforms in a few runs (red bars). In the conflict setting (Figure 3(a)), overall performance is lower and most G-Blending runs show larger drops than Uniform, although a small fraction still achieve marginal gains. Overall, these results suggest that while mitigating gradient conflicts can generally improve performance, the model can still underperform in certain cases.
Observing these results, we claim that gradient directions are essential for performance, but the reliability of gradients also has a significant impact on MML. A gradient can be viewed as a high-dimensional vector evolving throughout training, and each dimension is the partial derivative w.r.t. a model parameter (Goodfellow et al., 2016). We model gradients as random variables and describe them by probability distributions. We take two-dimensional updates as an example for further explanation (Figure 1). Uncertainty can be heterogeneous across dimensions and sources: the unimodal gradient (yellow) is more certain in the $x$-axis to update toward the negative $x$ direction, and the multi-modal gradient (blue) in the $y$-axis to update toward the negative $y$ direction. However, either by naive additive fusion or fusion after resolving conflicts, the obtained update differs from the more certain components of the two gradients, suggesting that uncertain gradient components may distort components where the other is reliable. Consequently, the update may be suboptimal or even harmful, leading to limited performance. This corresponds to the phenomenon we observed. In Section 5, the experiments suggest that taking into account uncertainty during optimization can bring further improvement in model performance. Thus, not only the gradient directions are important, but also the reliability of the gradients play a significant role for MML. More details and results are provided in Appendix G. Motivated by this insight, we propose an uncertainty-aware method for gradient aggregation for efficient optimization.
4 Methodology
Based on the above analyses, we propose BOGC-MML, a Bayesian-oriented gradient calibration method for MML. It calibrates the update direction during MML optimization by incorporating the uncertainty of each gradient. Specifically, we first model the gradient of each modality as a probability distribution using Bayesian inference. We then derive the gradient distribution for the multi-modal fusion loss (Section 4.1). Next, we propose an effective method to estimate evidence from these probability distributions, i.e., using a power mapping that is based on the precision of the probability distributions (Theorem 4.1). Obtaining the estimated evidence, we quantify the uncertainties of the modality-specific gradients and the gradient of the multi-modal fusion loss, drawing from subjective logic and evidence theory. Finally, based on the quantified uncertainties, we perform uncertainty-driven gradient aggregation using the reduced Dempster’s combination rule to obtain a more precise update direction, calibrating MML optimization (Section 4.2).
4.1 Bayesian-based Gradient Probability Distribution
In this subsection, we quantify modality-gradient uncertainty by inferring their probability distributions using Bayesian inference. We treat the parameters of the last pre-output layer as random variables and estimate their posteriors using a Laplace approximation (MacKay, 1992). Then, we express each modality gradient as a quadratic function of these parameters and approximate its distribution via first- and second-order moments. Finally, given per-modality gradient distributions, we derive the gradient distribution of the multi-modal fusion loss under Gaussian priors.
Posterior Estimation of Parameters. The network architectures of MML typically consist of modality-specific encoders, modality fusion layers, and classification heads (Baltrušaitis et al., 2018; Ngiam et al., 2011). We consider the parameters $\boldsymbol{\Theta}$ of the last layer before output as random variables and compute their probability distributions. Assuming that the posterior distribution of $\boldsymbol{\Theta}$ is a Gaussian distribution, which is a mild assumption (Neal, 2012), we approximate the posterior using the Laplace approximation (MacKay, 1992). Given the observed data $\mathcal{D}$, the posterior distribution can be formulated through Bayesian rule $p(\boldsymbol{\Theta}|\mathcal{D})\propto p(\mathcal{D}|\boldsymbol{\Theta})\cdot p(\boldsymbol{\Theta})$. Then, we approximate the posterior distribution of the parameters via a second-order Taylor expansion and Laplace approximation: for the $i$-th modality, the posterior distribution of $\boldsymbol{\Theta}^{i}$ is:
$\begin{aligned} \log p(\boldsymbol{\Theta}^{i}\mid\mathcal{D})&\approx\log p(\hat{\boldsymbol{\Theta}}^{i}\mid\mathcal{D})+(-\nabla_{\boldsymbol{\Theta}^{i}}\log p(\mathcal{D},\boldsymbol{\Theta}^{i}))^{\top}\Delta\boldsymbol{\Theta}^{i}\\ &+\frac{1}{2}(\Delta\boldsymbol{\Theta}^{i})^{\top}\mathbf{H}^{i}\Delta\boldsymbol{\Theta}^{i},\end{aligned}$
where $\hat{\boldsymbol{\Theta}}^{i}$ is the learned point estimate of the parameters, $\Delta\boldsymbol{\Theta}^{i}=\boldsymbol{\Theta}^{i}-\hat{\boldsymbol{\Theta}}^{i}$ is the deviation from the estimate, $\mathbf{q}=-\nabla_{\boldsymbol{\Theta}}\log p(\mathcal{D},\boldsymbol{\Theta})$, and $\mathbf{H}^{i}=-\nabla_{\boldsymbol{\Theta}^{i}}^{2}\log p(\mathcal{D},\boldsymbol{\Theta}^{i})$ is the negative Hessian matrix of the joint log-likelihood with respect to $p(\mathcal{D},\boldsymbol{\Theta}^{i})$. Then, we omit the first-order term (i.e. $(-\nabla_{\boldsymbol{\Theta}^{i}}\log p(\mathcal{D},\boldsymbol{\Theta}^{i}))^{\top}\Delta\boldsymbol{\Theta}^{i}$) in the Taylor expansion. As a result, the posterior takes the quadratic form: $\log p(\boldsymbol{\Theta}^{i}\mid\mathcal{D})\approx c^{i}+\frac{1}{2}(\Delta\boldsymbol{\Theta}^{i})^{\top}\mathbf{H}^{i}\Delta\boldsymbol{\Theta}^{i}$, where $c^{i}$ absorbs terms independent of $\boldsymbol{\Theta}^{i}$. The corresponding Gaussian distribution is (with details in Appendix C.3):
$p(\boldsymbol{\Theta}^{i}\mid\mathcal{D})\approx\mathcal{N}(\boldsymbol{\Theta}^{i}\mid(\hat{\boldsymbol{\Theta}}^{i}-(\mathbf{H}^{i})^{-1}\mathbf{q}^{i}),(\mathbf{H}^{i})^{-1}).$ (3)
However, the Hessian $\mathbf{H}^{i}$ is often not positive-definite (Achituve et al., 2024; Choromanska et al., 2015; Dauphin et al., 2014). We replace it with the generalized Gauss-Newton matrix, which provides a more stable approximation:
$\tilde{\mathbf{H}}^{i}=\sum_{j=1}^{|\mathcal{D}_{batch}|}(\mathbf{J}_{j}^{i})^{\top}\mathbf{B}_{j}^{i}\mathbf{J}_{j}^{i}+(\mathbf{V}_{p}^{i})^{-1},$ (4)
where $|\mathcal{D}_{batch}|$ is the batch size, $\mathbf{J}_{j}^{i}=\nabla_{\boldsymbol{\Theta}^{i}}\mathbf{f}^{i}(X_{j};\boldsymbol{\Theta}^{i})$ is the Jacobian of the model output for the $i$-th modality with respect to its parameters, $\mathbf{B}_{j}^{i}=\nabla_{\mathbf{f}^{i}}^{2}(-\log p(Y_{j}^{i}\mid X_{j},\boldsymbol{\Theta}^{i}))$ is the Hessian of the negative log-likelihood with respect to the model outputs, and $\mathbf{V}_{p}^{i}$ is the covariance of the Gaussian prior for $\boldsymbol{\Theta}^{i}$. Thus, the posterior approximation is:
$p(\boldsymbol{\Theta}^{i}\mid\mathcal{D})\approx\mathcal{N}(\boldsymbol{\Theta}^{i}\mid(\hat{\boldsymbol{\Theta}}^{i}-(\mathbf{H}^{i})^{-1}\mathbf{q}^{i}),\tilde{\mathbf{H}}^{i}).$ (5)
Gradient Distribution of Single Modality. Obtaining the posterior distributions, we then approximate the gradient probability distribution of the modalities. Specifically, we first establish the functional dependence of the gradient on the parameters. Next, we compute the first and second-order moments of the gradient through Monte Carlo sampling. Finally, the gradient distribution is approximated by a Gaussian distribution, from which the corresponding mean and covariance are derived.
Given the $i$-th modality and a random batch of samples $\mathcal{D}_{batch}\sim\mathcal{D}$, the gradient of the square loss with respect to the hidden layer $\mathbf{\psi}$ for the $t$-th sample is derived as:
$\mathbf{g}_{t}^{i}=\frac{\partial\mathcal{L}_{t}^{i}}{\partial\hat{Y}_{t}^{i}}\frac{\partial\hat{Y}_{t}^{i}}{\partial\mathbf{\psi}_{t}}=2\boldsymbol{\Theta}^{i}\mathbf{\psi}_{t}^{\top}\boldsymbol{\Theta}^{i}-2\boldsymbol{\Theta}^{i}Y_{t}^{i},$ (6)
where $\mathcal{L}_{t}^{i}=(Y^{i}-\hat{Y}^{i})^{2}$ is the square loss between label $Y^{i}$ and network output $\hat{Y}^{i}$ for modality i.
Eq. 6 defines the functional relationship between the gradient and the model parameters. Specifically, $\mathbf{g}_{t}^{i}$ is a quadratic function of the random variable $\boldsymbol{\Theta}^{i}$, which follows a Gaussian distribution. Consequently, $\mathbf{g}_{t}^{i}$ is also a random variable, and it follows a generalized chi-square distribution (Davies, 1973). The true distribution of $\mathbf{g}_{t}^{i}$ is generally intractable, as it does not admit a closed-form expression for its probability density (Bishop and Nasrabadi, 2006; Jordan et al., 1999). To enable efficient computation and inference, we approximate it with a Gaussian distribution (Rice and Rice, 2007). This choice is motivated by the fact that the Gaussian distribution has a closed-form density and is determined by its mean and covariance, making it analytically convenient (Bishop and Nasrabadi, 2006; Murphy, 2012). In practice, we estimate the parameters of this Gaussian approximation by matching the first- and second-order moments of $\mathbf{g}_{t}^{i}$, following the moment-matching principle (Minka, 2001).
In classification tasks, the gradient exhibits a nonlinear dependency on the model parameters, which precludes a closed-form solution for the first- and second-order moments (Bishop and Nasrabadi, 2006; Williams and Rasmussen, 2006). To address this, we first approximate the gradient distribution by assuming a Gaussian distribution, and then match its first- and second-order moments to make it as close as possible to the true distribution. Since directly obtaining these moments from the true distribution is intractable (Bishop and Nasrabadi, 2006; Welling and Teh, 2011), we use Monte Carlo sampling (Robert et al., 1999) to approximate them. Monte Carlo sampling only requires drawing samples from the posterior distribution and computing the gradient for each sample, without needing to rely on the specific structure of the model (Robert et al., 1999; Welling and Teh, 2011). Thus, given $n$ samples, we can derive the mean and covariance of the Gaussian approximation as follows:
$\begin{aligned} &\boldsymbol{\mu}_{t}^{i}:=\mathbb{E}[\mathbf{g}^{i}_{t}]\approx\frac{1}{n}\sum_{j=1}^{n}\mathbf{g}_{t}^{i}(\boldsymbol{\Theta}_{j}^{i}),\quad\\ &\boldsymbol{\Sigma}_{t}^{i}:=\mathbb{E}[(\mathbf{g}^{i}_{t})^{2}]-\boldsymbol{\mu}_{t}^{i}(\boldsymbol{\mu}_{t}^{i})^{\top}\approx\frac{1}{n}\sum_{j=1}^{n}\mathbf{g}_{t}^{i}(\boldsymbol{\Theta}_{j}^{i})\mathbf{g}_{t}^{i}(\boldsymbol{\Theta}_{j}^{i})^{\top}-\boldsymbol{\mu}_{t}^{i}(\boldsymbol{\mu}_{t}^{i})^{\top},\end{aligned}$
where $\boldsymbol{\Lambda}_{t}^{i}=\frac{1}{\boldsymbol{\Sigma}_{t}^{i}}$ is the precision matrix, $\boldsymbol{\Sigma}_{t}^{i}$ and $\boldsymbol{\Lambda}_{t}^{i}$ are both $K$-dimensional matrices, and $\boldsymbol{\Theta}_{j}^{i}$ are samples from $p(\boldsymbol{\Theta}^{i}|\mathcal{D}_{batch})$. The probability distribution of $\mathbf{g}_{t}^{i}$ is:
$p(\mathbf{g}_{t}^{i})\approx\mathcal{N}(\mathbf{g}_{t}^{i}\mid\boldsymbol{\mu}_{t}^{i},\boldsymbol{\Sigma}_{t}^{i}).$ (8)
Gradient Distribution of the Multi-Modal Fusion Loss. To obtain the gradient probability distribution of $\mathcal{L}_{\rm fusion}$, we proceed in two steps: (i) derive the expression of the gradient of $\mathcal{L}_{\rm fusion}$, and (ii) calculate the gradient distribution based on the Gaussian priors (Papoulis, 1965).
For the first step, in MML, the fusion loss $\mathcal{L}_{\rm fusion}$ can be expressed as $\mathcal{L}_{\rm fusion}=\mathcal{L}(\mathcal{C}(\zeta^{1}\oplus\zeta^{2}\oplus\cdots\oplus\zeta^{m}),y)$, where each modality is processed by a different modality-specific encoder $\zeta^{i}$ with parameter $\theta^{i}$ (Figure 2), $\mathcal{C}(\cdot)$ is a classifier, and $\oplus$ denotes a fusion operation (e.g., concatenation). Given the $t$-th sample, we obtain the the gradient of $\mathcal{L}_{\rm fusion}$, i.e., $\mathbf{g}_{t}^{\rm fusion}=\frac{1}{M}\sum_{i=1}^{M}\mathbf{g}_{t}^{i}$. Furthermore, since the Gaussian distribution remains Gaussian after weighted sum (Papoulis, 1965), we can obtain the mean and covariance of $\mathbf{g}_{t}^{\rm fusion}$ based on the gradient probability distributions of each modality, which can be expressed as:
[…]
where $\boldsymbol{\Sigma}_{t}^{\rm fusion}$ is also a $K$-dimensional matrix. Thus, the probability distribution of $\mathbf{g}_{t}^{\rm fusion}$ is:
$p(\mathbf{g}_{t}^{\rm fusion})\approx\mathcal{N}(\mathbf{g}_{t}^{\rm fusion}\mid\boldsymbol{\mu}_{t}^{\rm fusion},\boldsymbol{\Sigma}_{t}^{\rm fusion}).$ (10)
In summary, in this subsection, we first capture the uncertainty of each modality by modeling its gradient distribution through the posterior distribution. This approach yields a probabilistic description of the individual gradient spaces. Building on these per-modality distributions, we derive the Gaussian distribution for the fused gradient $\mathbf{g}_{t}^{\rm fusion}$ of the multi-modal fusion loss $\mathcal{L}_{\rm fusion}$.
Table 1: Classification accuracy(%) on CREMA-D and Kinetics Sounds. The best results are shown in bold. * indicates that the unimodal accuracy (Acc audio and Acc video) is obtained by fine-tuning a unimodal classifier with the trained unimodal encoder frozen.
Method CREMA-D Kinetics Sounds
Acc Acc audio Acc video Worst Acc Acc Acc audio Acc video Worst Acc
Audio-only - 61.69 - - - 53.63 - -
Video-only - - 56.05 - - - 49.20 -
Unimodal pre-trained & fine-tune 71.51 60.08 60.22 68.51 68.75 53.49 50.07 66.75
One joint loss* 66.13 59.27 36.56 63.63 64.61 52.03 35.47 61.61
Uniform baseline 71.10 63.44 51.34 69.30 68.31 53.20 40.55 66.11
G-Blending (Wang et al., 2020) 72.01 60.62 52.23 69.31 68.90 52.11 41.35 65.40
OGM (Peng et al., 2022)* 69.19 56.99 40.05 66.89 66.79 51.09 37.86 64.90
Greedy (Wu et al., 2022)* 67.61 60.69 38.17 65.51 65.32 50.58 35.97 62.52
PMR (Fan et al., 2023)* 66.32 59.95 32.53 64.62 65.70 52.47 34.52 62.30
AGM (Li et al., 2023)* 70.06 60.38 37.54 68.06 66.17 51.31 34.83 63.67
TMC (Han et al., 2022) 72.84 $\pm$ 0.98 61.73 $\pm$ 1.02 52.25 $\pm$ 0.88 69.76 $\pm$ 0.93 70.17 $\pm$ 1.12 55.53 $\pm$ 0.86 50.83 $\pm$ 1.21 67.64 $\pm$ 0.82
MMPareto (Wei and Hu, 2024) 75.13 65.46 55.24 71.93 70.13 56.40 53.05 67.33
D&R (Wei et al., 2025b) 73.52 61.94 52.59 70.13 69.10 54.18 44.86 66.75
ARL (Wei et al., 2025a) 74.19 62.15 52.91 70.50 68.40 53.10 40.70 66.01
ECML (Yue et al., 2025) 74.68 $\pm$ 1.17 62.31 $\pm$0.95 53.16 $\pm$ 0.83 70.77 $\pm$ 1.07 71.25 $\pm$ 1.17 56.24 $\pm$ 0.95 51.35 $\pm$ 1.17 68.08 $\pm$ 0.87
BOGC-MML 78.91 $\pm$ 0.78 68.52 $\pm$ 0.85 58.79 $\pm$ 0.59 74.02 $\pm$ 0.76 72.64 $\pm$ 0.82 58.11 $\pm$ 0.79 55.89 $\pm$ 0.91 69.37 $\pm$ 0.65
$\Delta$ (v.s. SOTA) $\uparrow$ 3.78 $\uparrow$ 3.06 $\downarrow$ 1.43 $\uparrow$ 2.09 $\uparrow$ 1.39 $\uparrow$ 1.71 $\uparrow$ 2.84 $\uparrow$ 1.29
4.2 Uncertainty Driven Gradient Aggregation
Based on the above results, we aggregate the gradient $\mathbf{g}_{t}^{i}$ of each modality with the gradient $\mathbf{g}_{t}^{\rm fusion}$ of the multi-modal loss, taking into account the uncertainty in each modality. The aggregated results are used to determine the optimal update direction. Specifically, inspired by subjective logic (Jsang, 2018) and evidence theory (Bao et al., 2021; Sensoy et al., 2018), we propose an effective method to convert each gradient into Dirichlet concentration parameters and associated evidence values. Obtaining these results, we compute the belief masses and uncertainty for each gradient in every dimension. Next, we fuse these beliefs and uncertainties across the gradients using the reduced Dempster’s combination rule, yielding a joint belief and a joint uncertainty per dimension. Finally, we weight and sum $\mathbf{g}_{t}^{i}$ and $\mathbf{g}_{t}^{\rm fusion}$ according to the fused belief and uncertainty, producing the calibrated update direction for the $i$-th modality.
Given the $i$-th modality and $t$-th sample, the covariance matrix $\boldsymbol{\Sigma}^{i}_{t}$ and the precision matrix $\boldsymbol{\Lambda}_{t}^{i}$ are both $K$-dimension (Eq.7). The element of the $d$-th row and $d$-th column in $\boldsymbol{\Lambda}_{t}^{i}$ is the precision of $\mathbf{g}_{t}^{i}$ in the $d$-th dimension. Similarly, the precision of $\mathbf{g}_{t}^{\rm fusion}$ in the $d$-th dimension can be obtained according to Eq.10, i.e., $\boldsymbol{\lambda}_{t,d}^{i}=(\boldsymbol{\Sigma}^{i}_{t})^{-1}[d,d]$ and $\boldsymbol{\lambda}_{t,d}^{\rm fusion}=(\boldsymbol{\Sigma}^{\rm fusion}_{t})^{-1}[d,d]$. Then, we propose:
Theorem 4.1.
Given a Gaussian distribution $\theta\sim\mathcal{N}\bigl(\mu_{0},\;\tfrac{1}{\lambda_{0}}\bigr)$. Given $n$ observations $x_{1},\dots,x_{n}$ and the observation noise precision (inverse variance) is $\tau>0$, the likelihood is $p(x_{1:n}|\theta)=\prod_{i=1}^{n}\mathcal{N}(x_{i}|\theta,\tfrac{1}{\tau})$. Given a real number exponent $s>0$ and a mapping $\phi\colon(0,\infty)\to(0,\infty)$ where $\phi(\lambda)=\lambda^{s}$, its scalar evidence can be estimated as $e=\phi(\lambda_{n})\;=\bigl(\lambda_{0}+n\tau\bigr)^{s}$.
Theorem 4.1 provides a method to convert the parameter posterior precision, which is obtained by combining a Gaussian prior with observed data into a single evidence value via a continuous monotonic mapping. The prior and data merge into a posterior precision (inverse variance), and then an exponent-controlled mapping translates that precision into evidence. The resulting evidence both captures the amount of information and allows us to adjust how sensitively it responds to changes in precision, unifying prior knowledge and observed uncertainty into one coherent measure.
Table 2: Test performance on CREMA-D and Kinetics Sounds where bold and underline represent the best and runner-up respectively. The network follows transformer-based framework MBT (Nagrani et al., 2021).
Method CREMA-D Kinetics Sounds
from scratch with pretrain from scratch with pretrain
Acc macro F1 Acc macro F1 Acc macro F1 Acc macro F1
One joint loss 44.96 42.78 66.69 67.26 42.51 41.56 68.30 69.31
Uniform baseline 45.30 43.74 69.89 70.11 43.31 43.08 69.40 69.60
G-Blending (Wang et al., 2020) 46.38 45.16 69.91 70.01 44.69 44.19 69.41 69.47
OGM-GE (Peng et al., 2022) 42.88 39.34 65.73 65.88 41.79 41.09 69.55 69.53
Greedy (Wu et al., 2022) 44.49 42.76 66.67 67.26 43.31 43.08 69.62 69.75
PMR (Fan et al., 2023) 44.76 42.95 65.59 66.07 43.75 43.21 69.67 69.87
AGM (Li et al., 2023) 45.36 43.81 66.54 67.75 43.65 43.57 69.59 69.14
TMC (Han et al., 2022) 46.83 $\pm$ 0.83 46.26 $\pm$ 0.93 68.96 $\pm$ 0.85 69.61 $\pm$ 1.02 43.91 $\pm$ 0.79 43.95 $\pm$ 0.84 69.02 $\pm$ 0.69 68.56 $\pm$ 0.72
ECML (Yue et al., 2025) 48.00 $\pm$ 0.78 47.26 $\pm$ 0.97 70.11 $\pm$ 0.69 70.86 $\pm$ 0.95 45.13 $\pm$ 0.74 44.75 $\pm$ 0.93 70.12 $\pm$ 0.82 69.61 $\pm$ 0.92
MMPareto (Wei and Hu, 2024) 48.66 48.17 70.43 71.17 45.20 45.26 70.28 70.11
BOGC-MML 52.21 $\pm$ 0.69 50.06 $\pm$ 0.74 74.00 $\pm$ 0.78 73.01 $\pm$ 0.83 47.19 $\pm$ 0.68 47.36 $\pm$ 0.86 73.17 $\pm$ 0.79 73.08 $\pm$ 0.83
$\Delta$ (v.s. SOTA) $\uparrow$ 3.55 $\uparrow$ 1.89 $\uparrow$ 3.57 $\uparrow$ 1.84 $\uparrow$ 1.99 $\uparrow$ 2.10 $\uparrow$ 2.89 $\uparrow$ 2.97
Table 3: Performance comparison when 50% samples are corrupted with Gaussian noise, i.e., zero mean with the variance of $N$. “(N, Avg.)” and “(N, Worst.)” denotes the average and worst-case accuracy. The best results are highlighted in bold.
Method RGB-Depth Scene Understanding Vision-Language Classification
NYU Depth V2 SUN RGB-D FOOD 101 MVSA
(0,Avg.) (0,Worst.) (10,Avg.) (10,Worst.) (0,Avg.) (0,Worst.) (10,Avg.) (10,Worst.) (0,Avg.) (0,Worst.) (10,Avg.) (10,Worst.) (0,Avg.) (0,Worst.) (10,Avg.) (10,Worst.)
CLIP (Sun et al., 2023) 69.32 68.29 51.67 48.54 56.24 54.73 35.65 32.76 85.24 84.20 52.12 49.31 62.48 61.22 31.64 28.27
ALIGN (Jia et al., 2021) 66.43 64.33 45.24 42.42 57.32 56.26 38.43 35.13 86.14 85.00 53.21 50.85 63.25 62.69 30.55 26.44
MaPLe (Khattak et al., 2023) 71.26 69.27 52.98 48.73 62.44 61.76 34.51 30.29 90.40 86.28 53.16 40.21 77.43 75.36 43.72 38.82
CoOp (Jia et al., 2022) 67.48 66.94 49.43 45.62 58.36 56.31 39.67 35.43 88.33 85.10 55.24 51.01 74.26 73.61 42.58 37.29
VPT (Jia et al., 2022) 62.16 61.21 41.05 37.81 54.72 53.92 33.48 29.81 83.89 82.00 51.44 49.01 65.87 64.98 32.79 29.21
Late fusion (Wang et al., 2016) 69.14 68.35 51.99 44.95 62.09 60.55 47.33 44.60 90.69 90.58 58.00 55.77 76.88 74.76 55.16 47.78
ConcatMML (Zhang et al., 2021) 70.30 69.42 53.20 47.71 61.90 61.19 45.64 42.95 89.43 88.79 56.02 54.33 75.42 75.33 53.42 50.47
AlignMML (Wang et al., 2016) 70.31 68.50 51.74 44.19 61.12 60.12 44.19 38.12 88.26 88.11 55.47 52.76 74.91 72.97 52.71 47.03
ConcatBow (Zhang et al., 2023) 49.64 48.66 31.43 29.87 41.25 40.54 26.76 24.27 70.77 70.68 35.68 34.92 64.09 62.04 45.40 40.95
ConcatBERT (Zhang et al., 2023) 70.56 69.83 44.52 43.29 59.76 58.92 45.85 41.76 88.20 87.81 49.86 47.79 65.59 64.74 46.12 41.81
MMTM (Joze et al., 2020) 71.04 70.18 52.28 46.18 61.72 60.94 46.03 44.28 89.75 89.43 57.91 54.98 74.24 73.55 54.63 49.72
TMC (Han et al., 2022) 71.06 69.57 53.36 49.23 60.68 60.31 45.66 41.60 89.86 89.80 61.37 61.10 74.88 71.10 60.36 53.37
LCKD (Wang et al., 2023a) 68.01 66.15 42.31 40.56 56.43 56.32 43.21 42.43 85.32 84.26 47.43 44.22 62.44 62.27 43.52 38.63
UniCODE (Xia et al., 2023) 70.12 68.74 44.78 42.79 59.21 58.55 46.32 42.21 88.39 87.21 51.28 47.95 66.97 65.94 48.34 42.95
SimMMDG (Dong et al., 2023) 71.34 70.29 45.67 44.83 60.54 60.31 47.86 45.79 89.57 88.43 52.55 50.31 67.08 66.35 49.52 44.01
MMBT (Kiela et al., 2019) 67.00 65.84 49.59 47.24 56.91 56.18 43.28 39.46 91.52 91.38 56.75 56.21 78.50 78.04 55.35 52.22
QMF (Zhang et al., 2023) 70.09 68.81 55.60 51.07 62.09 61.30 48.58 47.50 92.92 92.72 62.21 61.76 78.07 76.30 61.28 57.61
MNL (Gong et al., 2025) 71.52 68.75 56.72 52.15 62.25 60.45 49.10 47.80 93.03 91.45 63.16 60.50 78.85 76.80 61.82 57.90
BOGC-MML 74.62 72.15 58.31 54.72 65.48 64.21 50.37 47.93 94.20 93.10 64.10 62.85 79.60 78.90 62.90 59.40
$\Delta$ (v.s. SOTA) $\uparrow$ 3.10 $\uparrow$ 1.86 $\uparrow$ 1.51 $\uparrow$ 2.27 $\uparrow$ 3.04 $\uparrow$ 2.45 $\uparrow$ 1.27 $\uparrow$ 0.13 $\uparrow$ 1.05 $\uparrow$ 0.38 $\uparrow$ 0.94 $\uparrow$ 1.09 $\uparrow$ 0.75 $\uparrow$ 0.86 $\uparrow$ 1.08 $\uparrow$ 1.15
Table 4: Top-1 accuracy(%) under missing modalities on the BraTS Dataset. $\bullet$ and $\circ$ denote present and missing modalities, respectively.
Modalities Enhancing Tumour Tumour Core Whole Tumour
Fl T1 T1c T2 HMIS RSeg mmFm LCKD BOGC-MML HMIS RSeg mmFm LCKD BOGC-MML HMIS RSeg mmFm LCKD BOGC-MML
$\bullet$ $\circ$ $\circ$ $\circ$ 11.78 25.69 39.33 45.48 47.97 26.06 53.57 61.21 72.01 75.56 52.48 85.69 86.10 89.45 92.36
$\circ$ $\bullet$ $\circ$ $\circ$ 10.16 17.29 32.53 43.22 46.42 37.39 47.90 56.55 66.58 69.21 57.62 70.11 67.52 76.48 78.96
$\circ$ $\circ$ $\bullet$ $\circ$ 62.02 67.07 72.60 75.65 77.86 65.29 76.83 75.41 83.02 85.94 61.53 73.31 72.22 77.23 79.11
$\circ$ $\circ$ $\circ$ $\bullet$ 25.63 28.97 43.05 47.19 49.77 57.20 57.49 64.20 70.17 73.01 80.96 82.24 81.15 84.37 86.95
$\bullet$ $\bullet$ $\circ$ $\circ$ 10.71 32.13 42.96 48.30 51.21 41.12 60.68 65.91 74.58 77.32 64.62 88.24 87.06 89.97 93.03
$\bullet$ $\circ$ $\bullet$ $\circ$ 66.10 70.30 75.07 78.75 81.56 71.49 80.62 77.88 85.67 87.42 68.99 88.51 87.30 90.47 92.71
$\bullet$ $\circ$ $\circ$ $\bullet$ 30.22 33.84 47.52 49.01 53.11 57.68 61.16 69.75 75.41 78.25 82.95 88.28 87.59 90.39 93.78
$\circ$ $\bullet$ $\bullet$ $\circ$ 66.22 69.06 74.04 76.09 78.98 72.46 78.72 78.59 82.49 86.85 68.47 77.18 74.42 80.10 83.54
$\circ$ $\bullet$ $\circ$ $\bullet$ 32.39 32.01 44.99 50.09 52.87 60.92 62.19 69.42 72.75 76.02 82.41 84.78 82.20 86.05 88.77
$\circ$ $\circ$ $\bullet$ $\bullet$ 67.83 69.71 74.51 76.01 79.31 76.64 80.20 78.61 84.85 86.89 82.48 85.19 82.99 86.49 89.72
$\bullet$ $\bullet$ $\bullet$ $\circ$ 68.54 70.78 75.47 77.78 80.41 76.01 81.06 79.80 85.24 88.10 72.31 88.73 87.33 90.50 93.24
$\bullet$ $\bullet$ $\circ$ $\bullet$ 31.07 36.41 47.70 49.96 51.92 60.32 64.38 71.52 76.68 79.70 83.43 88.81 87.75 90.46 92.82
$\bullet$ $\circ$ $\bullet$ $\bullet$ 68.72 70.88 75.67 77.48 80.21 77.53 80.72 79.55 85.56 87.76 83.85 89.27 88.14 90.90 93.37
$\circ$ $\bullet$ $\bullet$ $\bullet$ 69.92 70.10 74.75 77.60 79.82 78.96 80.33 80.39 84.02 87.33 83.94 86.01 82.71 86.73 89.11
$\bullet$ $\bullet$ $\bullet$ $\bullet$ 70.24 71.13 77.61 79.33 83.01 79.48 80.86 85.78 85.31 88.61 84.74 89.45 89.64 90.84 93.76
According to Theorem 4.1, for the $d$-th dimension, we construct a power mapping $\phi(\cdot)$ to calculate the evidence $e_{t,d}^{i}$ of $\mathbf{g}_{t}^{i}$. Then, we compute Dirichlet concentration parameter $\alpha_{t,d}^{i}$ of $\mathbf{g}_{t}^{i}$. Also, we can obtain the evidence $e_{t,d}^{\rm fusion}$ and Dirichlet concentration parameter $\alpha_{t,d}^{\rm fusion}$ of $\mathbf{g}_{t}^{\rm fusion}$.
$e_{t,d}^{i}=\phi(\boldsymbol{\lambda}^{i}_{t,d})=(\boldsymbol{\lambda}^{i}_{t,d})^{s},$ $\quad\alpha_{t,d}^{i}=e_{t,d}^{i}+1,$ (11)
$e_{t,d}^{\rm fusion}=\phi(\boldsymbol{\lambda}_{t,d}^{\rm fusion})=(\boldsymbol{\lambda}^{\rm fusion}_{t,d})^{s},$ $\quad\alpha_{t,d}^{\rm fusion}=e_{t,d}^{\rm fusion}+1.$
Then, according to subjective logic (Jsang, 2018) and evidence theory (Bao et al., 2021; Sensoy et al., 2018), we compute the belief masses (i.e. $b_{t,d}^{i}$ and $b_{t,d}^{\rm fusion}$) and uncertainties (i.e. $u_{t}^{i}$ and $u_{t}^{\rm fusion}$) of $\mathbf{g}_{t}^{i}$ and $\mathbf{g}_{t}^{\rm fusion}$ respectively. It can be expressed as follows:
[…]
where $S_{t}^{i}$ and $S_{t}^{\rm fusion}$ are known as Dirichlet strength. After that, we construct sets of mass $\mathbb{M}^{i}_{t}=\{\{b_{t,d}^{i}\}_{d=1}^{K},u_{t}^{i}\}$ and $\mathbb{M}^{\rm fusion}_{t}=\{\{b_{t,d}^{\rm fusion}\}_{d=1}^{K},u_{t}^{\rm fusion}\}$ with the Dempster-Shafer rules, and combine them to obtain a joint mass $\mathbb{M}=\{b_{t,d},u_{t}\}$ according to Eq.27:
$b_{t,d}=\frac{b_{t,d}^{i}b_{t,d}^{\rm fusion}+b_{t,d}^{i}u_{t}^{\rm fusion}+b_{t,d}^{\rm fusion}u_{t}^{i}}{1-C},\quad u_{t}=\frac{u_{t}^{i}u_{t}^{\rm fusion}}{1-C}$
where $C=\sum_{p\neq q}b_{tp}^{i}b_{tq}^{\rm fusion}$ is a measure of the amount of conflict between the two mass sets. Then, we can obtain the gradient $\mathbf{g}_{\rm DS}$ after aggregation as follows:
$\mathbf{g}_{\rm DS}=\sum_{t=1}^{n}\sum_{d=1}^{K}b_{t,d}(b^{i}_{t,d}\boldsymbol{\mu}_{t,d}^{i}+b_{t,d}^{\rm fusion}\boldsymbol{\mu}_{t,d}^{\rm fusion}),$
where $\boldsymbol{\mu}_{t,d}^{i}$ and $\boldsymbol{\mu}_{t,d}^{\rm fusion}$ are the values of $\boldsymbol{\mu}_{t}^{i}$ and $\boldsymbol{\mu}_{t}^{\rm fusion}$ in the $d$-th dimension, respectively.
To this end, we can obtain a more effective and efficient gradient optimization direction considering the gradient uncertainty across modalities. Briefly, our method first captures the full uncertainty of each gradient source by modeling it as a probability distribution, yielding a richer representation of the gradient space. We then convert each gradient distribution’s precision into a scalar evidence measure, which in turn allows us to quantify the uncertainty of every gradient. Finally, we fuse these per-modality uncertainties using a reduced Dempster-Shafer rule to calibrate the combined update direction. Thereby, our method reduces excessive perturbations in sensitive dimensions and improves update efficiency in insensitive dimensions, leading to more accurate and efficient optimization in MML.
5 Experiment
In this section, we conduct a series of experiments to evaluate BOGC-MML on various benchmark datasets. More details and experiments are provided in Appendices D-G.
5.1 Experimental Settings
Datasets. We evaluate our method on seven datasets covering a range of multi-modal tasks: (i) audio-visual classification on CREMA-D (Cao et al., 2014) and Kinetics Sounds (Arandjelovic and Zisserman, 2017); (ii) RGB-Depth scene understanding on NYU Depth V2 (Silberman et al., 2012) and SUN RGB-D (Song et al., 2015); (iii) vision-language classification on UPMC FOOD101 (Wang et al., 2015) and MVSA (Niu et al., 2016); (iv) multi-modal medical segmentation on BraTS (Menze et al., 2014).
Implementation Details. Unless otherwise specified, all experiments are conducted using a ResNet-18 backbone as the default architecture configuration, with models trained from scratch without any pre-trained weights. We learn the posterior distribution over the parameters of the final layer in each modality-specific encoder. For optimization, we use stochastic gradient descent (SGD) with a momentum of 0.9 and a scaling factor $\gamma=1.5$. The initial learning rate is set to 0.1 and can be linearly scaled if necessary. All results are averaged over five runs on NVIDIA Tesla V100 GPUs.
5.2 Results
Performance Comparison. Following (Wei and Hu, 2024), we conduct experiments on CREMA-D and Kinetics Sounds datasets. We compare BOGC-MML with representative baselines, e.g., G-Blending (Wang et al., 2020), OGM-GE (Peng et al., 2022), Greedy (Wu et al., 2022), PMR (Fan et al., 2023), AGM (Li et al., 2023), TMC (Han et al., 2022), MMPareto (Wei and Hu, 2024), D&R (Wei et al., 2025b), ARL (Wei et al., 2025a), and ECML (Yue et al., 2025). Table 1 and 2 indicate the advantages of BOGC-MML.
Performance under Few-Shot and Zero-Shot Settings. To evaluate the effectiveness of different fusion strategies, we conduct experiments on four benchmark datasets (NYU Depth V2, SUN RGBD, FOOD101, and MVSA) under 0-shot and 10-shot scenarios. We compare a broad set of representative multi-modal baselines, such as LCKD (Wang et al., 2023a), UniCODE (Xia et al., 2023), SimMMDG (Dong et al., 2023), MMBT (Kiela et al., 2019), QMF (Zhang et al., 2023), and MNL (Gong et al., 2025). Table 3 further verify the effectiveness of BOGC-MML.
Performance on Missing Modality. To further evaluate the performance of BOGC-MML, we conduct comparative experiments under missing modality. We assess BOGC-MML against several strong baselines (Bakas et al., 2018; Wang et al., 2023a, 2025) across all 15 possible missing modality combinations on the BraTS dataset. Table 4 demonstrate the advantage of BOGC-MML in practice.
Figure 4: Results with uncertainty in non-conflict scenarios.
Figure 5: Parameter sensitivity about the parameter $s$.
Ablation Study about Parameter Sensitivity of Hyperparameter $s$. We conduct an experiment to evaluate the impact of $s$ on Kinetics-Sounds, with the range $[0.01,0.05,0.1,0.3,0.5,0.7,0.9]$. Figure 5 indicate that (i) when $s=0.5$ achieves the best performance, also our choice in this work; and (ii) the parameter $s$ has a relatively minor influence on the performance.
6 Conclusion
This paper rethinks MML from an optimization perspective. Our empirical analysis reveals that performance instability exists even in the absence of gradient conflicts. To account for this, we posit that gradient uncertainty is a critical yet overlooked factor, providing a new explanation for the interesting phenomenon. Building on this argument, we develop BOGC-MML, a framework that shifts the paradigm to explicitly modeling gradient uncertainty. Specifically, we treat gradients as random variables represented by probability distributions. We further design a mapping from gradient precision to evidence, enabling the use of a reduced Dempster’s rule to prioritize reliable information during aggregation. Extensive experiments on various benchmarks demonstrate the effectiveness and advantages of our approach.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning and Multi-Modal Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix
The appendix is organized into several sections:
  • Appendix A provides the detailed proofs of the theorem in the main text.
  • Appendix B provides details for all notations used in this paper.
  • Appendix C provides more discussion about the proposed methodology.
  • Appendix D provides details for the benchmark datasets used in the experiments.
  • Appendix E provides details for the baselines mentioned in the main text.
  • Appendix F contains details for the implementation of the experiments.
  • Appendix G provides the full results and analyses of the experiments.
Appendix A Proofs
Proof. Given a Gaussian prior $\theta\sim\mathcal{N}\bigl(\mu_{0},\tfrac{1}{\lambda_{0}}\bigr)$ with the mean $\theta$, where $\lambda_{0}>0$ is the precision. Given $n$ observations $x_{1},\dots,x_{n}$ with known noise precision $\tau>0$, the likelihood is $p(x_{1:n}\mid\theta)=\prod_{i=1}^{n}\mathcal{N}(x_{i}\mid\theta,\tfrac{1}{\tau})$. The prior density is $p(\theta)=\sqrt{\frac{\lambda_{0}}{2\pi}}{\rm exp}[-\frac{\lambda_{0}}{2}(\theta-\mu_{0})^{2}]$ and the likelihood factorizes as $p(x_{1:n}\mid\theta)=\prod_{i=1}^{n}\sqrt{\frac{\tau}{2\pi}}{\rm exp}[-\frac{\tau}{2}(x_{i}-\theta)^{2}]$. According to Bayes’ rule, we can obtain $p(\theta\mid x_{1:n})\propto p(\theta)p(x_{1:n}\mid\theta)$. Hence, we can derive
$p(\theta\mid x_{1:n})\propto{\rm exp}[-\frac{\lambda_{0}}{2}(\theta-\mu_{0})^{2}-\frac{\tau}{2}\sum_{i=1}^{n}(\theta-x_{i})^{2}]$ (15)
Then, we expand the sum of squares:
$\lambda_{0}(\theta-\mu_{0})^{2}+\tau\sum_{i=1}^{n}(\theta-x_{i})^{2}$ (16)
$=\lambda_{0}(\theta^{2}-2\mu_{0}\theta+\mu_{0}^{2})+\tau\sum_{i=1}^{n}(\theta^{2}-2x_{i}\theta+x_{i}^{2})$
$=(\lambda_{0}+n\tau)\theta^{2}-2(\lambda_{0}\mu_{0}+\tau{\textstyle\sum_{i}x_{i}})\theta+(\lambda_{0}\mu_{0}^{2}+\tau{\textstyle\sum_{i}x_{i}^{2}}),$
Define $\lambda_{n}=\lambda_{0}+n\tau,\quad\mu_{n}=\frac{\lambda_{0}\mu_{0}+\tau{\textstyle\sum_{i=1}^{n}x_{i}}}{\lambda_{n}}$. Then the exponent becomes
$\lambda_{n}(\theta^{2}-2\mu_{n}\theta+\mu_{n}^{2})+\underbrace{(\lambda_{0}\mu_{o}^{2}+\tau{\textstyle\sum_{i}x_{i}})-\lambda_{n}\mu_{n}^{2}}_{const.\;w.r.t\;\theta}$ (17)
Ignoring the additive constant, we can obtain $p(\theta\mid x_{1:n})\propto{\rm exp}[-\frac{\lambda_{n}}{2}(\theta-\mu_{n})^{2}]$. Thus, the posterior is Gaussian with updated precision $\lambda_{n}=\lambda_{0}+n\tau$ as follows:
$\theta\mid x_{1:n}\sim\mathcal{N}(\mu_{n},\frac{1}{\lambda_{n}}),$ (18)
where $\lambda_{n}$ objectively measures the total information accumulated from prior and $n$ data points.
Given an exponent $s\in\mathbb{R}^{+}$, we define a power mapping as follows:
$\phi:(0,\infty)\to(0,\infty),\quad\phi(\lambda)=(\lambda)^{s},$ (19)
where $\phi$ has three properties and we’ll verify them next.
First, $\phi$ has monotonicity, i.e. given any $\lambda_{2}>\lambda_{1}>0$, $\phi(\lambda_{2})>\phi(\lambda_{1})$. Since $s>0$, the function $\phi(\lambda)=(\lambda)^{s}$ has derivative
$\phi^{\prime}(\lambda)=s(\lambda)^{s-1},\quad\forall\lambda>0,$ (20)
A continuously differentiable function whose derivative is strictly positive is strictly increasing. Thus, $\lambda_{2}>\lambda_{1}\Rightarrow\lambda_{2}^{s}>\lambda_{1}^{s}$.
Second, $\phi$ control amplification or attenuation through $s$. Compute the second derivative:
$\phi^{\prime\prime}(\lambda)=s(s-1)\lambda^{s-2},$ (21)
If $s>1$, then $s(s-1)>0,\phi^{\prime\prime}(\lambda)>0$, so $\phi$ is convex. If $0<s<1$, then $s(s-1)<0,\phi^{\prime\prime}(\lambda)<0$, so $\phi$ is concave.
For convex $\phi$, the secant slope exceeds the tangent at the left endpoint:
$\frac{\phi(\lambda_{2})-\phi(\lambda_{1})}{\lambda_{2}-\lambda_{1}}>\phi^{\prime}(\lambda_{1})=s\lambda_{1}^{s-1},$ (22)
Multiplying both sides by $\lambda_{2}-\lambda_{1}>0$ yields $\phi(\lambda_{2})-\phi(\lambda_{1})>s\lambda_{1}^{s-1}(\lambda_{2}-\lambda_{1})$, so the differences in $\lambda$ are amplified when $s>1$.
For concave $\phi$, the reverse inequality holds $\phi(\lambda_{2})-\phi(\lambda_{1})<s\lambda_{1}^{s-1}(\lambda_{2}-\lambda_{1})$, so the differences in $\lambda$ are attenuated when $0<s<1$.
Third, we discuss the boundary behaviors of $\phi$. Specifically, as $\lambda\to 0^{+}$, $\lim_{\lambda\to 0^{+}}\lambda^{s}=0^{s}=0$. This interprets that with no information ($\lambda\to 0$), evidence $e\to 0$. As $\lambda\to+\infty$, $\lim_{\lambda\to+\infty}\lambda^{s}=+\infty$. This interprets that with unbounded information, evidence grows unbounded.
These three points establish that defining evidence by the power mapping is both mathematically valid (positive and strictly increasing) and semantically meaningful (treating precision as effective sample size and permitting controlled scaling).
Appendix B Notation
In this section, we briefly describe the symbols that we mainly use in this article. We give the definitions of notation according to their role.
  • $\square$
    Notations of Data
    • $\bullet$
      $\mathcal{X}$, $\mathcal{Y}$: the input space and label space.
    • $\bullet$
      $(x_{t},y_{t})\in\mathcal{X}\times\mathcal{Y}$: the $t$-th sample $x_{t}\in\mathcal{X}$ and the label $y_{t}\in\mathcal{Y}$.
    • $\bullet$
      $D^{i}=\{(x_{t}^{i},y_{t})\}_{t=1}^{n}$: the dataset of the $i$-th modality with $n$ samples.
    • $\bullet$
      $\mathcal{D}_{batch}\sim\mathcal{D}$: the random batch of samples.
    • $\bullet$
      $x_{t}=\{x_{t}^{1},\ldots,x_{t}^{(i)}\}$: each sample, which contains $M$ modalities.
    • $\bullet$
      $\Delta^{K-1}$: the $K$-dimensional unit simplex.
  • $\square$
    Notations of Model
    • $\bullet$
      $f_{\boldsymbol{\Theta}}\colon\mathcal{X}\to\mathcal{Y}$: the MML model.
    • $\bullet$
      $\mathbf{\psi}_{t}$: the hidden layer of MML models.
    • $\bullet$
      $\boldsymbol{\Theta}$: parameters of the last layer before output.
    • $\bullet$
      $\zeta^{i}$: the $i^{th}$ modality encoder.
    • $\bullet$
      $\theta^{i}$: parameters of the $i^{th}$ modality encoder.
  • $\square$
    Notations of Variables
    • $\bullet$
      $X,Y$: variables of the sample and corresponding label.
    • $\bullet$
      $\hat{Y}$: the answer distribution induced by the episode representation $u_{k,l}(x)$, where $\mathcal{A}$ is the answer space.
    • $\bullet$
      $M$: the total number of modalities.
    • $\bullet$
      $\boldsymbol{\mu},\boldsymbol{\Sigma}$: the mean and covariance matrix of gradient distribution.
    • $\bullet$
      $\boldsymbol{\mu}_{t}^{i},\boldsymbol{\Sigma}_{t}^{i}$: the mean and covariance matrix of gradient distribution $\mathbf{g}_{t}^{i}$ of the $i$-th modality for the $t$-th sample.
    • $\bullet$
      $(\boldsymbol{\sigma}_{t}^{i})^{2}$: the variance of the covariance matrix $\boldsymbol{\Sigma}_{t}^{i}$.
    • $\bullet$
      $\boldsymbol{\Lambda}_{t}^{i}:=\frac{1}{\boldsymbol{\Sigma}_{t}^{i}}$: the precision of gradient distribution $\mathbf{g}_{t}^{i}$ of the $i$-th modality for the $t$-th sample.
    • $\bullet$
      $\boldsymbol{\lambda}_{t,d}^{i}$: the precision of the gradient distribution $\mathbf{g}_{t}^{i}$ of the $i$-th modality for the $t$-th sample in the $d$-th dimension.
    • $\bullet$
      $(\boldsymbol{\Sigma}^{i}_{t})^{-1}[d,d]$: the element of the $d$-th row and $d$-th column in $\boldsymbol{\Lambda}_{t}^{i}$.
    • $\bullet$
      $e^{i}_{t,d},b^{i}_{t,d}$: the evidence and belief of the gradient distribution $\mathbf{g}_{t}^{i}$ of the $i$-th modality for the $t$-th sample in the $d$-th dimension.
    • $\bullet$
      $u_{t}^{i}$: the uncertainty of gradient distribution $\mathbf{g}_{t}^{i}$ of the $i$-th modality for the $t$-th sample.
    • $\bullet$
      $\boldsymbol{\alpha}$: the concentration parameters of Dirichlet distribution.
    • $\bullet$
      $\mathbb{M}=\{b,u\}$: the set of masses including belief and uncertainty.
    • $\bullet$
      $\boldsymbol{S}$: the Dirichlet strength.
    • $\bullet$
      $\boldsymbol{C}$: measure of the amount of conflict between the two mass sets.
  • $\square$
    Notations of Learning Objective
    • $\bullet$
      $\mathcal{L}_{\rm fusion}$: the multi-modal fusion loss that aggregates information from all modalities.
    • $\bullet$
      $\mathcal{L}_{\rm uni}^{i}$: the loss for the individual $i$-th modality.
    • $\bullet$
      $\mathcal{L}=(Y-\hat{Y})^{2}$: the standard square loss between label $Y$ and network output $\hat{Y}$.
    • $\bullet$
      $\mathbf{g}_{t}^{i}$: the gradient distribution of the $i$-th modality for the $t$-th sample.
    • $\bullet$
      $\mathbf{g}_{t}^{\rm fusion}$: the gradient distribution of the multi-modal fusion loss.
    • $\bullet$
      $\mathbf{g}_{\rm DS}$: the gradient update direction obtained by BOGC-MML.
Appendix C More Discussion
C.1 Illustration of Important Concepts and Theories
C.1.1 Bayesian Inference
Bayesian inference represents and manages uncertainty by constructing probabilistic models. In Bayesian inference (Box and Tiao, 2011; Dempster, 1968), uncertainty is described using probability distributions that reflect the degree of belief in different possible values of a variable or parameter. Bayesian inference provides a principled foundation for reasoning under uncertainty, and has been widely adopted in machine learning to quantify model confidence (Box and Tiao, 2011; Dempster, 1968). The key idea is to treat model parameters as random variables and to update our belief about them after observing data. This update is formalized by Bayes’ theorem
$p(\boldsymbol{\Theta}\mid\mathcal{D})=\frac{p(\mathcal{D}\mid\boldsymbol{\Theta})p(\boldsymbol{\Theta})}{p(\mathcal{D})}$ (23)
where $p(\boldsymbol{\Theta})$ is the prior distribution encoding our initial belief about the parameters $\boldsymbol{\Theta}$, $p(\mathcal{D}\mid\boldsymbol{\Theta})$ is the likelihood of observing the data given the parameters, and $p(\boldsymbol{\Theta}\mid\mathcal{D})$ is the resulting posterior after incorporating evidence from the data.
C.1.2 Dirichlet Distribution
The Dirichlet distribution is a continuous probability distribution defined on a $K$-dimensional unit simplex $\Delta^{K-1}=\{\boldsymbol{z}=(z_{1},\cdots,z_{K})\mid z_{n}>0,\sum_{n=1}^{K}z_{n}=1\}$ (Bishop and Nasrabadi, 2006). The probability density function of the Dirichlet distribution is as follows:
$Dir(\boldsymbol{z}\mid\boldsymbol{\alpha})=\frac{1}{{\rm B}(\boldsymbol{\alpha})}\prod_{n=1}^{K}z_{n}^{\alpha_{n}-1}$ (24)
where $\boldsymbol{\alpha}=[\alpha_{1},\cdots,\alpha_{K}]$ are concentration parameters of the Dirichlet distribution, ${\rm B}(\boldsymbol{\alpha})$ is the $K$-dimensional multinomial beta function (Johnson et al., 1972).
In multi-modal classification, given a multi-modal observation $\boldsymbol{x}=\{x^{i}\}^{M}_{i=1},\boldsymbol{x}\in\mathcal{X}$ and the label $\boldsymbol{y}\in\mathcal{Y}$, where $\mathcal{X}$ is the set of all observations and $\mathcal{Y}$ is the set of all labels. Given the $i$-th modality, the generative process is as follows:
$\boldsymbol{z}^{i}\sim Dir(\boldsymbol{z}^{i}\mid\boldsymbol{\alpha}^{i}),\quad\boldsymbol{y}\sim Mult(\boldsymbol{y}\mid\boldsymbol{z}^{i})$ (25)
where $Mult(\boldsymbol{y}\mid\boldsymbol{z}^{i})$ is a multinomial distribution and $\boldsymbol{z}^{i}=[z^{i}_{1},\cdots,z^{i}_{1}]$ are the parameters of the multinomial distribution.
C.1.3 Subjective Logic and Evidence Theory
Subjective logic (Jsang, 2018) and evidence theory (Bao et al., 2021; Sensoy et al., 2018) offer a theoretical framework for modeling uncertainty using the Dirichlet distribution (Appendix C.1.2). Specifically, in the context of multi-modal $K$-class classification, for the $i$-th modality, the Dirichlet distribution $Dir(\boldsymbol{\mu}^{i}\mid\boldsymbol{\alpha}^{i})$ is used, where the parameters $\boldsymbol{\alpha}^{i}=[\alpha^{i}_{1},\dots,\alpha^{i}_{K}]$ are associated with several key concepts from subjective logic and evidence theory, including the evidence $\boldsymbol{e}^{i}=[e^{i}_{1},\dots,e^{i}_{K}]$, belief $\boldsymbol{b}^{i}=[b^{i}_{1},\dots,b^{i}_{K}]$, and uncertainty $u^{i}$. The evidence refers to the metrics derived from the input that can support classification (Han et al., 2022). The relationship between the evidence $e^{i}_{k}$ and the Dirichlet parameter $\alpha^{i}_{k}$ is given by:
[…]
where $u^{i}\geq 0$ is the overall uncertainty for the $i$-th modality, $b^{i}_{k}\geq 0$ is the belief mass of $k$-th class, $S^{i}=\sum_{k=1}^{K}\alpha^{i}_{k}=\sum_{k=1}^{K}(e^{i}_{k}+1)$ is known as the Dirichlet strength. From Eq. 26, we can deduce that the more evidence $e^{i}_{k}$ is acquired for the $k$-th class, the higher the belief $b^{i}_{k}$ assigned to that class. Conversely, when less evidence $\boldsymbol{e}^{i}$ is obtained, the uncertainty $\boldsymbol{u}^{i}$ increases. Building on these concepts, in this work, we model modality-gradient uncertainty to calibrate optimization toward better update directions, weighting gradient updates by the reliability of each modality’s evidence.
C.1.4 Dempster-Shafer Evidence Theory
Dempster-Shafer evidence theory (DST) provides a mathematical framework for modeling and integrating uncertain or imprecise information (Dempster, 2008). A fundamental component of DST is Dempster’s combination rule, which allows for the fusion of evidence from multiple independent sources. To address the computational challenges of this operation, the Reduced Dempster’s combination rule was introduced as an efficient approximation method (Han et al., 2022). In the context of multi-modal $K$-class classification, given the $i$-th and $j$-th modalities, we construct two independent sets, i.e., $\mathbb{M}^{i}=\{\{b^{i}_{k}\}_{k=1}^{K},u^{i}\}$ and $\mathbb{M}^{j}=\{\{b^{j}_{k}\}_{k=1}^{K},u^{j}\}$, which are based on the beliefs and uncertainties of the $i$-th and $j$-th modalities, respectively. The combined set $\mathbb{M}$ is obtained by applying the combination rule $\mathbb{M}=\mathbb{M}^{i}\oplus\mathbb{M}^{j}$. The combination rule is defined as:
$b_{k}=\frac{b^{i}_{k}b^{j}_{k}+b^{i}_{k}u^{j}+b^{j}_{k}u^{i}}{1-C},\quad u=\frac{u^{i}u^{j}}{1-C}$
where $C=\sum_{p\neq q}b^{i}_{p}b^{j}_{q}$ is a measure of the amount of conflict between the two mass sets, $\frac{1}{1-C}$ is a scale factor for normalization. This combination process can be extended to incorporate more than two modalities. Assume we have $M$ modalities, the belief and gradient uncertainty for each modality can be combined iteratively using $\mathbb{M}=\mathbb{M}^{1}\oplus\mathbb{M}^{2}\cdots\oplus\mathbb{M}^{M}$. In this work, we propose a novel method for uncertainty aggregation of different gradients, inspired by the reduced Dempster’s combination rule, to more effectively guide the optimization process.
C.2 More Illustration of Figure 2
Figure 2a illustrates the prevailing loss-fusion paradigm in MML: (i) each modality is supervised by its own unimodal loss: $\mathcal{L}_{\text{uni}}^{i}$ and $\mathcal{L}_{\text{uni}}^{j}$; (ii) a multi-modal fusion loss $\mathcal{L}_{\text{fusion}}$ is introduced as a joint objective that simultaneously optimizes both modalities; (iii) the overall training objective is obtained via a weighted summation of unimodal and multi-modal losses. Figure 2b juxtaposes two kinds of gradient aggregation strategies: (i) deterministic gradient aggregation, where gradients are combined via fixed or dynamically learned scalar weights or multi-objective optimization to find a compromise direction, and (ii) uncertainty driven gradient aggregation, which first estimates uncertainty of each gradient and then utilizes these uncertainties to guide the MML optimization. Figure 2c provides the widely used MML optimization framework. Specifically, the input of each modality is passed through its dedicated encoder and merged into a hidden representation. Then, both the fusion loss and individual modality losses are computed. During backpropagation, we obtain both individual modality gradients and the standard fusion gradient. These gradients are merged into an update direction via a specific aggregation strategy. Finally, the aggregated gradient is applied through the optimizer (e.g. SGD or Adam) to update the parameters of MML network.
Compared with the deterministic gradient aggregation strategy, the optimal direction of the uncertainty driven gradient aggregation strategy is located at the highest density point of the two gradient distributions. This means that the optimal direction corresponds to the most probable gradient direction under both modality distributions simultaneously, ensuring the update step is the direction that best explains all available evidence. If a gradient source exhibits fluctuations or outliers in specific dimensions, the uncertainty driven gradient aggregation strategy naturally shifts weight toward the other gradient with more reliable estimates, attenuating the influence of noisy signals.
C.3 Derivations of Posterior Approximation
In this subsection, we provide the derivation of Eq.3. First, we define and recap some notations here.
$p=\log p(\boldsymbol{\Theta}\mid\mathcal{D}),\mathbf{q}=-\nabla_{\boldsymbol{\Theta}}\log p(\mathcal{D},\boldsymbol{\Theta}),\mathbf{H}=-\nabla_{\boldsymbol{\Theta}}^{2}\log p(\mathcal{D},\boldsymbol{\Theta})$ (28)
According to Eq.2, we use the above notations to obtain the following form:
[…]
$+\frac{1}{2}(\boldsymbol{\Theta}-(\hat{\boldsymbol{\Theta}}-\mathbf{H}^{-1}\mathbf{q}))^{\top}\mathbf{H}(\boldsymbol{\Theta}-(\hat{\boldsymbol{\Theta}}-\mathbf{H}^{-1}\mathbf{q}))$
The above takes the quadratic form of a Gaussian having mean $(\hat{\boldsymbol{\Theta}}-\mathbf{H}^{-1}\mathbf{q})$ and covariance $\mathbf{H}^{-1}$.
C.4 Explanation of Steps within BOGC-MML
In this subsection, we provide more explanations about our proposed method.
[…]
Why is the first-order term of Eq.2 omitted. In Eq.2, we expand the log-posterior around the learned point estimate $\hat{\boldsymbol{\Theta}}^{i}$. The first-order term $(-\nabla_{\boldsymbol{\Theta}^{i}}\log p(\mathcal{D},\boldsymbol{\Theta}^{i}))^{\top}\Delta\boldsymbol{\Theta}^{i}$ is omitted because the gradient of the joint log-density vanishes, i.e. $\nabla_{\boldsymbol{\Theta}^{i}}\log p(\mathcal{D},\boldsymbol{\Theta}^{i})|_{\boldsymbol{\Theta}^{i}=\hat{\boldsymbol{\Theta}}^{i}}=0$. Thus, when expanding around a stationary point, the Jacobian term is identically zero and contributes nothing to the approximation.
How can the gradient distribution of each modality be calculated through the first and second-order moments. By matching the first and second-order moments, we can obtain the Gaussian distribution that, among all Gaussian distributions, is closest in the sense of Kullback-Leibler divergence to the true distribution of the gradients (Minka, 2013). In other words, we preserve as much of the original distribution shape as possible by forcing it into Gaussian form.
Why $\alpha$ equals $e+1$. According to subjective logic (Jsang, 2018) and evidence theory (Bao et al., 2021; Sensoy et al., 2018), each Dirichlet concentration parameter $\alpha$ represents a “pseudo-count”, while the actual observed evidence is denoted $e$. To ensure a non-informative prior when no evidence is available, the initial pseudo-count is usually selected as 1 (Jsang, 2018). Consequently, we can obtain $\alpha=e+1$.
C.5 Related Work
C.5.1 Multi-Modal Learning
Multi-modal learning (MML) aims to integrate complementary signals from modalities such as images, text, and audio to improve performance (Baltrušaitis et al., 2018; Ngiam et al., 2011). A line of work tackles fusion from an optimization viewpoint by aggregating modality gradients and adaptively adjusting each modality’s contribution during training (Peng et al., 2022; Tsai et al., 2019; Wang et al., 2024, 2020). For example, (Wei and Hu, 2024) resolves conflicts between unimodal and multi-modal objectives via Pareto-based gradient integration; (Yue et al., 2025) uses an evidential dissonance measure to assess view quality and performs dissonance-aware belief integration with an evidential gradient penalty for robust multi-view fusion; (Jiang et al., 2025) modifies gradients via a flat gradient modification strategy (with SAM-style smoothing) to enhance cross-modal interaction and alleviate modality imbalance. However, these methods often rely on fixed coefficients and implicitly treat gradients as equally reliable, overlooking gradient uncertainty and dimension-wise sensitivity. We address this by explicitly modeling each modality gradient as a distribution and steering fusion toward more reliable update directions. Additional comparisons are provided in Appendix C.
C.5.2 Bayesian Theory
Bayesian theory provides a principled framework for modeling under uncertainty by treating model parameters as random variables and updating their probability distributions after observing data (Dempster, 1968; Box and Tiao, 2011). Specifically, given the observed data $\mathcal{D}$ and model parameters $\theta$, Bayesian theory defines how a prior belief $p(\theta)$ is revised into a posterior distribution $p(\theta|\mathcal{D})$ through the likelihood term $p(\mathcal{D}|\theta)$, i.e., $p(\theta|\mathcal{D})=\frac{p(\mathcal{D}|\theta)\,p(\theta)}{p(\mathcal{D})}$. Bayesian inference has been widely employed in deep learning to represent gradient uncertainty through parameter distributions. Bayes by Backprop (Blundell et al., 2015) approximates the posterior distribution over network weights and thus captures their uncertainty during optimization. Similarly, Bayesian convolutional networks (Gal and Ghahramani, 2015) model the uncertainty of convolutional filters, enabling confidence-aware prediction. More generally, Bayesian neural networks represent parameters as probability distributions rather than fixed values (Blundell et al., 2015; Gal and Ghahramani, 2016). These methods focus on uncertainty at the parameter level and thus model gradient uncertainty. In this work, our approach also builds on Bayesian principles, but focuses on modeling the probability distribution of modality gradients. Furthermore, by incorporating subjective logic, we transform the quantified uncertainty into belief scores, which provide an interpretable measure of confidence for each gradient dimension.
C.5.3 Subjective Logic and Evidence Theory
Subjective logic and evidence theory provide important theoretical foundations for modeling uncertainty under incomplete information (Jsang, 2018; Bao et al., 2021; Sensoy et al., 2018). Subjective logic (Jsang, 2018) extends traditional probabilistic logic by explicitly incorporating degrees of belief and uncertainty, enabling more nuanced representations of uncertain evidence. In parallel, evidence theory, also known as Dempster-Shafer theory (DST) (Dempster, 1968, 2008), offers a framework for combining uncertain information from multiple sources through evidence masses and combination rules. In the context of deep learning, Dirichlet-based evidence networks (Sensoy et al., 2018) have been proposed to model prediction uncertainty by generating evidence parameters that map to class-wise belief masses and uncertainty scores. Classical Dempster’s combination rule provides a principled way to merge beliefs, while reduced or adaptive variants have been developed to alleviate conflicts between modalities (Han et al., 2022; Zhang et al., 2023), making evidence theory more applicable in complex multi-modal environments.
In this work, we adopt a Bayesian-oriented perspective to directly model the probability distribution of modality gradients, and then employ subjective logic to transform uncertainty into dimension-wise belief scores. These belief scores are subsequently aggregated using a reduced evidence combination rule to calibrate the gradient update direction. In this way, we bridge Bayesian inference with subjective logic and evidence theory, moving beyond prediction calibration to explicitly guide optimization in MML.
C.6 More Comparison
In this subsection, we further discuss the differences between BOGC-MML and the previous or concurrent methods.
Recently, optimization methods using a combination of multi-modal fusion loss and multiple unimodal losses have achieved promising performance (Wang et al., 2020). Many methods focus on how to obtain multi-modal fusion losses through new strategies or how to aggregate multi-modal fusion losses with unimodal losses to obtain updated directions (Peng et al., 2022; Sener and Koltun, 2018; Tsai et al., 2019; Wang et al., 2024, 2020; Wei and Hu, 2024). However, these methods usually regard the lost gradient as deterministic values. Unlike previous MML methods, our proposed BOGC-MML presents a novel optimization framework. Specifically, we focus on how to model the probability distributions of multi-modal fusion loss and unimodal losses, and then how to use the probability distributions for gradient aggregation to obtain the updated direction.
For uncertainty-based learning, although the existing neural networks are effective in many fields, most neural networks rely on deterministic functions, so the uncertainties in neural networks are difficult to obtain. To address these uncertainties, several methods have been developed. Recently, methods based on subjective logic (Jsang, 2018) and Dempster-Shafer evidence theory (Dempster, 2008) have shown promise in uncertainty estimation. Unlike methods indirectly model uncertainty through network weights, these methods have been successively proposed to directly model uncertainty (Charpentier et al., 2020; Kopetzki et al., 2021; Malinin and Gales, 2018; Malinin et al., 2019; Sensoy et al., 2018). Notably, (Han et al., 2022) and (Yue et al., 2025) successfully pioneer the quantification of modality reliability. By deriving confidence scores from data evidence, they effectively assess the trustworthiness of different modalities. This allows the model to dynamically reweight the contribution of each modality, ensuring that the learning process prioritizes more reliable information sources.
Despite recent efforts to improve MML model performance, these approaches typically consider gradient sources as deterministic values and struggle to quantify the uncertainty of gradients. To this end, we propose a novel calibration framework BOGC-MML that quantifies and aggregate uncertainties of gradients to calibrate the update direction in MML optimization. We also compare our method with the previous uncertainty-focused methods, where experimental results demonstrate our advantages.
C.7 Broader Impacts and Limitations
In this subsection, we briefly illustrate the broader impacts and limitations of this work.
Broader Impacts. This work introduce BOGC-MML, a Bayesian-oriented gradient calibration framework that explicitly models and aggregates the uncertainty of modality-specific gradients and the gradient of the multi-modal fusion loss during MML optimization. Our empirical results demonstrate that the proposed method achieve validity across diverse benchmarks. This work bridges Bayesian inference, evidence theory, and gradient-based optimization, offering a blueprint for future work on uncertainty-aware training regimes across machine learning disciplines.
Limitations. This work evaluates BOGC-MML mainly on supervised tasks where loss gradients are driven by explicit ground-truth labels. However, we have not explored how uncertainty-calibrated gradient aggregation behaves in unsupervised or self-supervised settings, such as clustering, generative modeling, or contrastive representation learning without labels. We will investigate additional case studies to extend this work in the future.
Appendix D Benchmark Datasets
In this section, we provide a brief overview of the benchmark datasets used in our experiments.
  • CREMA-D (Cao et al., 2014) is a multi-modal emotion recognition dataset designed to study how humans perceive emotions through audio, visual, and audiovisual stimuli. It contains 7,442 clips from 91 actors (48 male, 43 female) of diverse racial and ethnic backgrounds (African American, Asian, Caucasian, Hispanic, etc.). Actors performed 12 sentences in six basic emotions (Anger, Disgust, Fear, Happy, Neutral, Sad) across four intensity levels (Low, Medium, High, Unspecified). Each clip was rated by 2,443 crowd-sourced participants in three modalities: audio-only, video-only, and audiovisual. The dataset includes annotations for emotion categories, intensity levels, and actor demographics (age, sex, race, ethnicity).
  • Kinetics Sounds (Arandjelovic and Zisserman, 2017) is a subset of the Kinetics-600 dataset, curated for audio-visual action recognition. It focuses on human actions with distinctive sound cues, such as playing musical instruments, cooking, or sports. The dataset includes  19,000 video clips (10 seconds each) across 31 action classes, derived from YouTube. Each clip is annotated with temporal and semantic labels for both audio and visual modalities. It is widely used to benchmark models that fuse audio and visual features for tasks like action classification and event detection.
  • NYU Depth V2 (Silberman et al., 2012) contains more than 1,400 indoor scene video sequences and over 400,000 paired RGB and depth images. It covers a diverse range of indoor environments such as bedrooms, kitchens, living rooms, and offices. The dataset is widely used for semantic segmentation and scene classification tasks. In multi-modal learning, it provides complementary RGB and depth modalities, enabling more robust indoor scene understanding.
  • SUN RGB-D (Song et al., 2015) is a large-scale benchmark for indoor scene understanding, consisting of approximately 10,000 RGB-D images annotated with over 800 object and scene categories. Compared to NYU Depth V2, it offers richer category diversity and higher scene complexity. SUN RGB-D is extensively used for scene recognition and 3D spatial reasoning, where the integration of RGB and depth modalities provides strong complementarity for multi-modal models.
  • UPMC FOOD101 (Wang et al., 2015) is a multi-modal benchmark designed for food recognition, comprising around 100,000 images across 101 food categories, each paired with corresponding recipe or textual descriptions. The dataset targets cross-modal classification tasks, where the goal is to jointly leverage image and text information for food category prediction. FOOD101 poses challenges due to the semantic gap between visual content and textual descriptions, as well as noise and ambiguity in the text modality.
  • MVSA (Niu et al., 2016) is a multi-modal sentiment analysis corpus, consisting of 4,869 image-text pairs collected from social media platforms. Each pair is annotated with sentiment polarity labels: positive, negative, or neutral. The dataset is primarily used to evaluate multi-modal models for sentiment classification, where images and text often provide complementary yet weakly aligned cues. MVSA is challenging because it requires models to capture subtle cross-modal semantic consistency under noisy and subjective contexts.
  • BraTS (Menze et al., 2014) is a multi-modal medical imaging dataset for brain tumor segmentation. While not explicitly detailed in the provided search results, it is widely recognized in medical AI research. The dataset includes multiparametric MRI scans (T1, T1c, T2, FLAIR) from glioblastoma patients, annotated with tumor sub-regions (enhancing tumor, tumor core, whole tumor). It is used in annual challenges to benchmark segmentation algorithms under missing or incomplete modalities.
Appendix E Baselines
In this section, we provide a brief overview of the baselines used for comparison in our experiments.
  • G-Blending (Wang et al., 2020) proposes gradient blending to dynamically adjust modality-specific gradients during training. It mitigates modality competition by reweighting gradients based on their contributions, improving worst-case accuracy.
  • Occlusion-guided Modulation (OGM) (Peng et al., 2022) adaptively modulates modality-specific learning speeds based on their relative contributions. It suppresses over-confident modalities to encourage balanced learning. While improving robustness, its unimodal performance drops significantly due to suppressed modality-specific features.
  • Greedy (Wu et al., 2022) propose a modality selection strategy that greedily optimizes the primary modality while neglecting the weaker one. This leads to severe video performance degradation, highlighting the risk of modality collapse in imbalanced settings.
  • Probability Margin Regularization (PMR) (Fan et al., 2023) enforces a margin between correct and incorrect class probabilities to reduce modality overconfidence. However, its overly strict constraints harm video modality learning, suggesting limited adaptability to severe imbalance.
  • Adaptive Gradient Modulation (AGM) (Li et al., 2023) dynamically adjusts gradient magnitudes for each modality based on their training status. While moderately improving worst-case accuracy, it struggles to balance modality-specific and joint representation learning.
  • Trusted Modality Combination (TMC) (Han et al., 2022) introduces a trusted modality combination (TMC) mechanism that first estimates a reliability score for each modality via a data-driven trust model, then weights modality gradients according to these scores during fusion.
  • Evidential Conflictive Multi-view Learning (ECML) (Yue et al., 2025) leverages evidential calibration to convert each modality predictive uncertainty into Dempster-Shafer evidence; these evidence masses are used to derive fusion weights that reflect both confidence and uncertainty.
  • MMPareto (Wei and Hu, 2024) is a multi-objective optimization framework that seeks Pareto-optimal solutions for modality balance. It achieves strong results by optimizing accuracy and fairness simultaneously. Its Pareto-efficient training ensures no modality is severely neglected.
  • D&R (Wei et al., 2025b) proposes a two-stage framework consisting of diagnosing and re-learning. It first diagnoses the learning status and contribution of each modality to identify imbalance. Subsequently, it employs a re-learning strategy to adjust the optimization process, ensuring that under-optimized modalities are effectively trained to achieve balanced multimodal learning.
  • ARL (Wei et al., 2025a) addresses the modality imbalance issue by improving multimodal learning via an imbalanced learning perspective. It introduces mechanisms to dynamically balance the contribution of different modalities, preventing the model from relying solely on the dominant modality while suppressing the weaker one.
  • CLIP (Sun et al., 2023) learns a shared embedding space by pretraining on large-scale paired image–text data. During inference, classification is performed via cosine similarity between image and text embeddings. As a strong vision–language foundation model, CLIP is often used as a baseline for multi-modal robustness and transferability evaluation.
  • ALIGN (Jia et al., 2021) is another large-scale vision–language pretraining framework that utilizes contrastive learning to align visual and textual representations. Compared to CLIP, ALIGN uses a larger training corpus and enables more generalizable cross-modal features, which we use as a prompt‐based baseline in our evaluation.
  • MaPLe (Khattak et al., 2023) introduces multi-modal prompt learning to adapt pretrained vision–language models to downstream tasks. It learns task‐dependent prompt tokens for each modality, effectively bridging the gap between pretraining objectives and downstream classification.
  • CoOp (Jia et al., 2022) proposes a context optimization strategy that learns continuous prompts in the embedding space of the text branch. In multi-modal tasks, CoOp is typically used to adapt pretrained text encoders to specific visual categories through prompt tuning.
  • VPT (Jia et al., 2022) focuses on visual prompt tuning by inserting a small set of learnable prompt tokens into the image encoder. This enables efficient adaptation of pretrained vision transformers without modifying the backbone parameters, and serves as a lightweight multi-modal tuning baseline.
  • Late fusion (Wang et al., 2016) concatenates unimodal outputs at the prediction (logit) level and performs weighting or averaging to obtain the final decision. Despite its simplicity, it is widely used as a baseline to evaluate whether sophisticated fusion strategies offer practical advantages.
  • ConcatMML (Zhang et al., 2021) concatenates representations from different modalities at the feature level and feeds them into a shared classifier. It provides a straightforward way to jointly utilize multi-modal information and is commonly adopted as a baseline for feature-level multi-modal fusion.
  • AlignMML (Wang et al., 2016) extends feature concatenation by introducing an alignment module to reduce modality discrepancies before fusion. By learning modality alignment in the latent space, AlignMML improves the effectiveness of joint representations and enhances downstream performance.
  • ConcatBow (Zhang et al., 2023) concatenates visual features with bag-of-words (BoW) textual representations to form a joint feature vector, which is then fed into a shared classifier. Unlike ConcatBERT or AlignMML that use transformer-based encoders, ConcatBow relies on lightweight BoW features to perform fusion and serves as a simple yet effective baseline for feature-level multi-modal integration.
  • ConcatBERT (Zhang et al., 2023) concatenates visual and textual features and encodes them using a pre-trained BERT model to produce multi-modal representations. Unlike simple concatenation, it leverages a transformer encoder to capture cross-modal interactions and therefore serves as a stronger feature-level fusion baseline.
  • MMTM (Joze et al., 2020) introduces a multi-modal transfer module that modulates features across modalities through channel-wise attention. It automatically adjusts the contribution of each modality based on feature importance and has been shown to improve robustness in audio–visual tasks.
  • TMC (Han et al., 2020) proposes a trusted region mechanism to regularize multi-modal fusion by suppressing unreliable modality features. It improves model robustness by dynamically weighting each modality and selectively fusing only trustworthy feature dimensions.
  • LCKD (Wang et al., 2023a) exploits latent cross-modal knowledge distillation to align modality-specific representations during training. By encouraging consistency between different modalities in the latent space, it improves robustness and ensures that gradient signals remain complementary rather than conflicting.
  • UniCODE (Xia et al., 2023) learns a unified codebook shared across modalities to represent multi-modal data in a modality-agnostic manner. This unified representation facilitates cross-modal transfer and fusion while reducing sensitivity to noisy inputs.
  • SimMMDG (Dong et al., 2023) constructs a multi-modal embedding space in which the gradient similarity between modalities is explicitly optimized. This design reduces gradient discrepancies during training and results in more stable and noise-resilient multi-modal fusion.
  • MMBT (Kiela et al., 2019) integrates multiple modality-specific branches and combines their predictions through learned mixture weights. It allows the model to emphasize different modalities depending on the input conditions and has been adopted in various multi-modal classification benchmarks.
  • QMF (Zhang et al., 2023) employs a quasi mixture-of-experts fusion strategy to dynamically combine unimodal and multi-modal losses. It provides flexibility in balancing the contributions of different modalities and often yields stable performance under perturbations such as noise or missing modalities.
  • MNL (Gong et al., 2025) introduces the concept of negative learning into multimodal classification. By explicitly training the model to reject incorrect classes across modalities (Multimodal Negative Learning), it reduces the risk of overfitting to noisy correlations and enhances the robustness of the joint representation.
In addition, for the missing modality scenario, we select several strong baseline as comparison baselines.
  • Hierarchical Multi-modal Image Segmentation (HMIS) (Havaei et al., 2016) is a traditional or early deep learning-based method for multi-modal medical image segmentation. It typically employs a hierarchical architecture to process multi-modal inputs sequentially or in parallel. When modalities are missing, HMIS may rely on simple heuristics (e.g., interpolation, modality replacement) or statistical imputation to fill gaps. However, its performance degrades significantly under extreme missing-modality scenarios due to limited adaptability. It serves as a foundational baseline for comparing robustness against modality incompleteness.
  • Recurrent Segmentation Network (RSeg) (Chen et al., 2019) leverages recurrent neural networks (RNNs) or attention-based recurrent mechanisms to iteratively refine segmentation predictions. It is designed to handle sequential dependencies between modalities or spatial regions. For missing modalities, RSeg might use learned memory cells or attention gates to compensate for absent data. While effective in capturing temporal/spatial relationships, its performance depends heavily on the quality of modality-specific feature extraction and may struggle with highly heterogeneous missing patterns.
  • Multi-modal Feature Memory (mmFm) (Zhang et al., 2022) introduces a memory-augmented network that stores modality-specific features in a shared memory bank. During inference, the model retrieves and combines features from the memory to compensate for missing modalities. This approach explicitly models cross-modal correlations and leverages stored patterns to handle incomplete inputs. However, its effectiveness relies on the diversity of the training data and may fail when encountering unseen missing-modality combinations.
  • Latent Cross-Modal Knowledge Distillation (LCKD) (Wang et al., 2023a) employs knowledge distillation to transfer knowledge from a teacher model (trained on complete modalities) to a student model that handles missing modalities. The latent representations of the teacher guide the student to mimic robust features even when specific modalities are absent. This method mitigates performance drops by enforcing consistency in latent space but requires a two-stage training process and may suffer from information loss during distillation.
Appendix F Implementation Details
Unless otherwise specified, our experiments utilize a ResNet-18 backbone, and all models are trained from scratch without pre-trained weights. We learn the parameters of the last layer as part of the posterior approximation. Moving on to the optimization process, we use SGD with momentum (0.9) and $\gamma=1.5$ in experiments. The initial learning rate is set to 0.1 and can be linearly scaled if needed. We also determine the hyperparameter $s$ through grid search, resulting in $s=0.5$. We have constructed dedicated ablation experiments for the selection of these parameters. Experimental results demonstrate that the model maintains relatively stable performance across different datasets under various hyperparameter settings. For evaluation, the training dataset is randomly split into training and validation sets. Hyperparameters are tuned on the validation set to maximize performance, and final accuracy is reported on the test dataset. All experiments are conducted over five runs using NVIDIA Tesla V100 GPUs.
Appendix G Full Experimental Results and Analyses
G.1 Performance under Few-Shot and Zero-Shot Settings
To evaluate the effectiveness of different fusion strategies, we conduct experiments on four benchmark datasets (NYU Depth V2, SUN RGBD, FOOD101, and MVSA) under both 0-shot and 10-shot scenarios. Both average accuracy and worst-case accuracy are reported to provide a view of model performance across different domains. The compared methods include a broad set of multi-modal baselines spanning from classical fusion strategies to recent advanced vision-language models, as summarized in Table 5 and Table 6.
Specifically, the baselines can be roughly grouped into five categories. (i) pretrained contrastive models such as CLIP (Sun et al., 2023) and ALIGN (Jia et al., 2021) leverage large-scale image–text pretraining and are widely used as strong backbones in multi-modal research. (ii) prompt learning methods like MaPLe (Khattak et al., 2023), CoOp (Jia et al., 2022), and VPT (Jia et al., 2022) adapt pretrained models by optimizing textual or visual prompts, aiming to improve generalization with minimal finetuning. (iii) fusion strategies include late fusion (Wang et al., 2016), Concat-based approaches (ConcatMML (Zhang et al., 2021), AlignMML (Wang et al., 2016), ConcatBow (Zhang et al., 2023), ConcatBERT (Zhang et al., 2023)), joint modules such as MMTM (Joze et al., 2020) and TMC (Han et al., 2020), and knowledge distillation-based LCKD (Wang et al., 2023a), all of which aim to integrate or calibrate visual and textual representations for improved multi-modal alignment. (iv) distribution generalization and robustness-oriented methods, including SimMMDG (Dong et al., 2023), UniCODE (Xia et al., 2023), and QMF (Zhang et al., 2023), focus on handling distribution shifts or providing theoretical robustness guarantees. (v) transformer-based multi-modal encoders such as MMBT (Kiela et al., 2019) offer another family of strong baselines by directly integrating visual and textual streams via joint attention layers.
As shown in Table 5 and Table 6, our method demonstrates superior robustness across four datasets (NYU Depth V2, SUN RGBD, FOOD101, and MVSA). Specifically, BOGC-MML achieves both higher average accuracy and worst-case accuracy compared to competitive baselines, and consistently maintains strong performance even when the noise variance increases. These results further verify the effectiveness of modeling gradient uncertainty and demonstrate that our method can provide more reliable gradient updates even under adverse training environments.
Table 5: Performance comparison of RGB-Depth scene understanding when 50% samples are corrupted with Gaussian noise, i.e., zero mean with the variance of $N$. “(N, Avg.)” and “(N, Worst.)” denotes the average and worst-case accuracy. The best results are highlighted in bold.
Method NYU Depth V2 SUN RGB-D
(0,Avg.) (0,Worst.) (10,Avg.) (10,Worst.) (0,Avg.) (0,Worst.) (10,Avg.) (10,Worst.)
[…]
$\Delta$ (v.s. SOTA) $\uparrow$ 3.10 $\uparrow$ 1.86 $\uparrow$ 1.51 $\uparrow$ 2.27 $\uparrow$ 3.04 $\uparrow$ 2.45 $\uparrow$ 1.27 $\uparrow$ 0.13
Table 6: Performance comparison of vision-language classification when 50% samples are corrupted with Gaussian noise, i.e., zero mean with the variance of $N$. “(N, Avg.)” and “(N, Worst.)” denotes the average and worst-case accuracy. The best results are highlighted in bold.
Method FOOD 101 MVSA
(0,Avg.) (0,Worst.) (10,Avg.) (10,Worst.) (0,Avg.) (0,Worst.) (10,Avg.) (10,Worst.)
[…]
$\Delta$ (v.s. SOTA) $\uparrow$ 1.05 $\uparrow$ 0.38 $\uparrow$ 0.94 $\uparrow$ 1.09 $\uparrow$ 0.75 $\uparrow$ 0.86 $\uparrow$ 1.08 $\uparrow$ 1.15
Figure 6: Trade-off performance of various methods, with point size reflecting training time. (a) Result on CREMA–D. (b) Result on Kinetics Sounds
Table 7: Effect of belief-guided representation. This experiment aims to examine whether the dimension-wise belief values learned by BOGC-MML provide meaningful guidance for optimization. “Original” denotes the feature representation directly extracted from the model. “Reduce” means we reduce the values of high-belief dimensions in the representation, while “Increase” denotes that the values of low-belief dimensions are amplified. All modified representations are subsequently normalized and fed into the frozen classifier for evaluation.
Datasets Metric BOGC-MML
Original Reduce Increase
CREMA-D (Cao et al., 2014) Accuracy(%) 78.91 77.45 77.91
Kinetics Sounds (Arandjelovic and Zisserman, 2017) Accuracy(%) 72.64 71.02 71.45
NYU Depth V2 (Silberman et al., 2012) (0,Avg) 74.62 73.20 73.71
SUN RGB-D (Song et al., 2015) (0,Avg) 65.48 63.85 64.42
FOOD 101 (Wang et al., 2015) (0,Avg) 94.20 92.48 93.16
MVSA (Niu et al., 2016) (0,Avg) 79.60 78.04 78.62
BraTS (Menze et al., 2014) Accuracy(%) 83.01 81.45 81.94
Table 8: Performance under noisy conditions on NYU Depth V2.
Method MMBT QMF BOGC-MML
Accuracy(%) 62.89 64.72 70.15
G.2 Performance under Noisy Conditions
We further examine the performance of BOGC-MML when the input data is corrupted. As shown in Section G.1, our method can narrow the performance gap between average and worst-case scenarios. A similar trend is observed when facing more severe conditions, such as missing modalities (Table 4).
To provide a more intuitive verification, we conduct an additional experiment following the setup in Section G.1. Specifically, we use NYU Depth V2 and apply a random mask to roughly 30% of the data area. We compare the performance of MMBT, QMF, and BOGC-MML. As summarized in Table 8, the baseline models experience a clear degradation under corruption, whereas BOGC-MML leads to a slight reduction. These findings indicate that BOGC-MML can maintain stable performance even when inputs are noisy or partially missing, demonstrating its robustness.
Figure 7: Impact of Posterior Modeling Depth on Accuracy and Training Cost.
Figure 8: Accuracy under different batch sizes.
Figure 9: BOGC-MML with different loss functions on the CREMA-D dataset.
G.3 Posterior Modeling Depth
Our proposed method learns the posterior distribution over the parameters of the final layer in each modality-specific encoder. To further investigate the impact of posterior modeling depth on model performance and training cost, we conduct a comparative study by varying the number of layers included in the posterior modeling. Specifically, we evaluate three configurations where the posterior distribution is learned over the last one, two, or three layers of each encoder. All experiments are conducted on the CREMA-DOG dataset using a standard CNN-based architecture. As shown in Figure 9, model performance remains largely consistent across different modeling depths, whereas training cost increases exponentially with the number of layers involved in posterior modeling.
G.4 Ablation Study about Batch Size
We investigate the impact of the training batch size. The values are chosen from $[32,64,128,256]$ with grid search. The experiment is conducted on the CREMA-D dataset, and the results are presented in Figure 9. From the results, we can observe that the batch size has little effect on our method, which proves the stability of the proposed method.
G.5 Ablation Study about Loss Functions
To further investigate the generalization capability of BOGC-MML, we analyze its performance under different loss function objectives. While our main framework derives gradient distributions based on Square Loss (MSE), we empirically evaluate the method using Cross-Entropy, KL Divergence, and Wasserstein distance on the CREMA-D dataset.
Figure 9 presents the trade-off between classification accuracy and training cost (measured in hours). We observe that BOGC-MML maintains performance superiority, demonstrating its robustness to the choice of optimization objective.
G.6 Trade-off Performance
We further investigate the trade-off between model performance and computational efficiency. Under the same backbone, we compare several baseline methods, including OGM, TMC, MMPareto, and BOGC, on the CREMA-D dataset. As shown in Figure 6, the vertical axis represents classification accuracy (%), while the horizontal axis indicates the evaluated methods. The size of each point is proportional to its corresponding training time, reflecting the computational cost. The results demonstrate that BOGC achieves a desirable balance between accuracy and training overhead, delivering both strong performance and efficiency.
Figure 10: Visualization examples of gradient conflict vs. gradient non-conflict between unimodal and multi-modal objectives. (a) non-conflict scenario: unimodal and multi-modal gradients are broadly aligned, (b) conflict scenario: unimodal and multi-modal gradient contain opposing components. The heatmaps visualize cosine similarity between the two gradients, where values >> 0 indicate acute angles (non-conflicting) and values << 0 indicate obtuse angles (conflicting).
G.7 Qualitative Analysis of Gradient Conflict
To further clarify the distinction between non-conflicting and conflicting gradients in multi-modal training, we construct the visualization using samples from CREMA-D (Cao et al., 2014) (Figure 10). Concretely, we draw some samples and compute the unimodal gradient and the multimodal gradient. Then, we quantify their cosine similarity. The cosine similarity >> 0 indicates acute angles and aligned (non-conflicting) updates, while the cosine similarity << 0 indicates obtuse angles and opposing (conflicting) directions.
We contrast two fusion outcomes: a naive fused update $\rm g_{ns}$ and a conflict-mitigated update $\rm g_{c}$. In the non-conflict regime, the dominant directions of ($\rm g_{\text{uni}}$) and ($\rm g_{\text{multi}}$) are aligned, so ($\rm g_{\text{ns}}$) typically follows their shared descent tendency. In the conflict regime, negative-similarity regions imply that naive fusion directly mixes opposing components, causing ($\rm g_{\text{ns}}$) to drift in an inefficient compromise direction. By suppressing or avoiding these opposing components, the corrected update ($\rm g_{c}$) remains closer to a compatible descent direction, which is more likely to translate into stable and effective optimization in the dataset.
Figure 11: Twelve visualization examples on the BraTS dataset. We employ the model to compute the dimensional belief values of randomly sampled data (resized to $32\times 32$). In the heat map, larger values reflect dimensions that contribute more strongly to the gradient fusion and are regarded as more reliable. (a)-(l) represent the belief map of these twelve samples.
G.8 Qualitative Analysis of Gradient Uncertainty
To further analyze the behavior of our uncertainty-aware gradient fusion, we also provide a qualitative visualization of the per-dimension belief values learned by BOGC-MML. In particular, we take the BraTS dataset as a case study and randomly select twelve training samples. For each sample, we compute the dimension-wise belief values $b_{t,d}$ over a resized $32\times 32$ grid. The resulting block maps are shown in Figure 11, where warm colors indicate dimensions with higher belief values, i.e., those that the model considers more reliable during gradient aggregation.
Across the twelve examples, we observe that the belief maps exhibit heterogeneous spatial patterns, highlighting different sets of informative dimensions depending on the input sample. In particular, some samples present a few scattered regions with relatively high belief, whereas others show multiple small clusters. Moreover, experiments in Section 5 indicate that optimizing based on beliefs and uncertainties can lead to better model performance. Thus, we suggest that the model adaptively identifies the dimensions that contribute more to the update direction, instead of assigning uniform weights across all gradients.
G.9 Practical Guidance of Belief Representation
To further investigate the role of belief in guiding model representations, we conduct a set of experiments that aims to examine whether the dimension-wise belief values in BOGC-MML provide meaningful guidance for representation learning. We conduct a set of experiments across multiple datasets, and the results are shown in Table 7. In the BraTS, we adopt the same four modalities, namely F1, T1, T1c, and T2.
Specifically, the “Original” setting denotes the unmodified representation directly extracted from the model. In the “Reduce” setting, we lower the weights of high-belief dimensions in the representation, which are expected to carry more informative features. Feeding these representations into the classifier leads to a notable drop in accuracy. Conversely, in the “Increase” setting, we amplify the values of low-belief dimensions by doubling them and then apply normalization, which also results in degraded performance. These results collectively indicate that the model’s belief assignment is not arbitrary but indeed captures the contribution of different representation dimensions.
These results indicate that the belief assignment learned by BOGC-MML is not arbitrary but indeed captures the contribution of different representation dimensions. When left unaltered, belief-weighted representations preserve the most informative structure and yield the best accuracy. In contrast, artificially reducing or inflating belief dimensions disrupts this balance. Importantly, the consistency of these findings across diverse datasets–spanning audio, vision, text, and medical imaging–demonstrates that the learned belief values serve as reliable and interpretable guidance for optimization.